### v1

# Image Generation from Contextually-Contradictory Prompts

Abstract.
Text-to-image diffusion models excel at generating high-quality, diverse images from natural language prompts. However, they often fail to produce semantically accurate results when the prompt contains concept combinations that contradict their learned priors. We define this failure mode as contextual contradiction, where one concept implicitly negates another due to entangled associations learned during training. To address this, we propose a stage-aware prompt decomposition framework that guides the denoising process using a sequence of proxy prompts. Each proxy prompt is constructed to match the semantic content expected to emerge at a specific stage of denoising, while ensuring contextual coherence. To construct these proxy prompts, we leverage a large language model (LLM) to analyze the target prompt, identify contradictions, and generate alternative expressions that preserve the original intent while resolving contextual conflicts. By aligning prompt information with the denoising progression, our method enables fine-grained semantic control and accurate image generation in the presence of contextual contradictions. Experiments across a variety of challenging prompts show substantial improvements in alignment to the textual prompt.
copyright: nonedoi: XXXXXXX.XXXXXXXsubmissionid: 1201
Figure 1. Our method addresses contextual contradictions in text-to-image generation. These contradictions arise when one concept implicitly conflicts with another due to the model’s learned associations. For example, if a concept like “nest” is strongly entangled with “birds”, it may conflict with another concept in the prompt such as “rooster”, leading the model to ignore or distort part of the semantic meaning.
1. Introduction
Text-to-image diffusion models have demonstrated remarkable capabilities in generating high-quality and diverse visual content from natural language prompts (Rombach et al., 2022; Ramesh et al., 2021; Ho et al., 2020). However, achieving precise semantic alignment between the generated image and the conditioning prompt remains an open challenge. This issue becomes particularly pronounced when the target prompt lies outside the model’s training distribution, such as prompts that combine semantically plausible yet statistically uncommon or unconventional concepts. For example, as shown in Figure 1, generating an image from the prompt “a rooster in a nest” often results in colorful birds sitting in a nest. This is due to the model’s strong prior entanglement between bird species and nests, which implicitly contradicts the notion of a rooster.
We refer to this phenomenon as Contextual Contradiction, a conflict between two concepts that arises not from direct semantic opposition, but from the model’s associations learned during training. More precisely, we say that concept $A$ contextually contradicts concept $B$ if the model’s prior entangles $A$ with concept $C$, and $B$ contradicts $C$. In Figure 1, we illustrate this with a blowing dragon, which stands in contextual contradiction with the water due to its entanglement with fire.
The phenomenon of contextual contradiction in text-to-image models relates to the broader issue of spurious correlations in deep learning. Models often rely on shortcuts, which are predictive patterns in the training data that do not reflect true semantics (Geirhos et al., 2020). In NLP, models exploit syntactic heuristics rather than reasoning (McCoy et al., 2019), while in computer vision, they may depend on background cues rather than object features (Xiao et al., 2021; Singh et al., 2020). For example, classifiers may fail to recognize cows on beaches due to background biases (Beery et al., 2018). In this paper, we identify similar biases in generative models, where contextual contradictions arise when prompts include plausible but uncommon concept combinations.
To address this issue, we build on the observation that the denoising process follows a coarse-to-fine progression, during which different semantic attributes (e.g., background, pose, shape, and texture) emerge at distinct stages (Chefer et al., 2023; Balaji et al., 2022; Patashnik et al., 2023). Our key idea is to guide the model at each stage of denoising with the information most relevant to the type of content being formed at that point. To achieve this, we decompose the original prompt into a sequence of proxy prompts, each aligned with the attributes expected at a specific stage and designed to avoid contextual contradictions.
Ensuring that proxy prompts preserve the original intent while avoiding contextual contradictions requires a broad understanding of the real world. For example, it involves understanding that a bear is entangled with specific poses, such as walking on all fours or standing upright, which in turn, contradicts the handstand pose. Our method leverages a large language model (LLM) to analyze the target prompt, identify contextual contradictions, and construct suitable proxy prompts.
We demonstrate that, by using in-context examples and prompting the LLM to follow a reasoning process through a brief explanation, it can identify contextually contradictory concepts in a prompt and determine the appropriate stage of denoising at which each attribute should be introduced. It can also suggest alternative, non-conflicting concepts that preserve the intended attributes and use them to construct stage-specific proxy prompts. In doing so, the LLM effectively guides the model toward the intended meaning of the original prompt.
Through extensive experiments, we demonstrate the effectiveness of our method in generating images from contextually contradicting text prompts. By introducing prompt information at targeted stages, our approach generates precise combination of semantic attributes while avoiding undesired entanglement of unrelated attributes. Compared to previous methods, our stage-dependent prompt decomposition leads to more faithful and semantically aligned generations.
2. Related Work
Learned Spurious Correlations
Machine learning models are known to be sensitive to spurious correlations in their training data (Geirhos et al., 2020; McCoy et al., 2019; Ye et al., 2024). This sensitivity often leads to a significant drop in performance when training-time associations between concepts do not hold at test time. Prior work has extensively studied this issue in discriminative computer vision tasks, showing, for example, that recognition models tend to rely heavily on background cues (Singh et al., 2020; Xiao et al., 2021; Beery et al., 2018).
In our work, we show that text-to-image models exhibit similar behavior. When given contextually contradictory prompts – combinations of concepts that conflict with correlations seen during training – diffusion models often fail to generate images that accurately reflect the prompt. We evaluate this behavior using the Whoops! dataset (Bitton-Guetta et al., 2023), which contains prompts constructed by first describing two co-occurring elements, and then replacing one with a less compatible alternative. This results in scenarios that are unlikely to occur in the real world.
Semantic Alignment in Text-to-Image Synthesis
Text-to-image models often struggle to fully capture the semantic intent of input prompts, particularly when prompts involve complex or internally conflicting concepts. Previous works have analyzed common failure cases and proposed targeted improvements across various stages of the generation pipeline, including enhanced text embedding representations (Rassin et al., 2022; Feng et al., 2022; Tunanyan et al., 2023), refined attention mechanisms (Dahary et al., 2024) and guidance strategies that leverage attention maps for loss heuristics (Chefer et al., 2023; Rassin et al., 2023; Meral et al., 2024; Agarwal et al., 2023). Despite these advances, existing methods often fail to handle prompts containing contradictory concepts arising from the model’s learned associations. Our work directly addresses this underexplored challenge, focusing on contextually contradictory prompts.
Multi-Prompt Conditioning Techniques
Conditioning diffusion models on multiple prompts has emerged as an effective strategy for improving control and compositionality. One line of work, primarily focused on personalization, introduces distinct learned tokens at different layers of the model and at various denoising timesteps (Alaluf et al., 2023; Voynov et al., 2023). This design allows each token to capture different attributes of the personalized concept, leading to improved identity preservation. Other approaches vary the prompt across timesteps to modulate specific visual properties, such as object shape (Liew et al., 2022; Patashnik et al., 2023), or alternate between rare and frequent object descriptions to improve attribute binding (Park et al., 2024). Fine-grained spatial control has also been achieved by assigning sub-prompts to separate image regions (Yang et al., 2024). Additionally, some methods leverage multiple diffusion models, each conditioned on different prompt attributes, and combine their outputs into a unified prediction (Liu et al., 2022; Bar-Tal et al., 2023). Unlike most prior works, we focus on utilizing multiple prompt to settle internal semantic tension, where concept combinations lead to contextual contradictions.
Figure 2. Our method navigates the denoising process using time-dependent proxy prompts. (Top) A large language model (LLM) is employed to decompose a target prompt with entangled or conflicting attributes into a sequence of proxy prompts and their corresponding time-step intervals. (Bottom) These intermediate prompts are injected into the diffusion process at the specified intervals. This time-dependent prompt conditioning aligns more closely with the model’s prior at each stage, enabling a semantically consistent and temporally coherent progression through the denoising process.
[…]
LLM-Guided Diffusion
LLMs have demonstrated strong capabilities in language understanding. Moreover, they possess a broad, implicit grasp of world knowledge, acquired through large-scale training on diverse textual data. Recent approaches have leveraged these capabilities to guide diffusion model generation, often incorporating planning and reasoning to improve semantic alignment (Yang et al., 2024; Park et al., 2024; Hu et al., 2024). In our work, we employ LLMs with in-context learning to identify contextual contradictions, generate proxy prompts, and determine the corresponding timestep ranges for conditioning, while promoting a reasoning process through brief, explanatory outputs.
3. Method
In this section, we begin by analyzing the coarse-to-fine behavior of the denoising process (Section 3.1). Building on the insights from our analysis, we introduce our training-free framework for resolving contextual contradictions in text-to-image generation. As illustrated in Figure 2, our approach consists of two main components: (i) prompt decomposition (Section 3.2, top part of the figure), and (ii) stage-aware prompt injection (Section 3.3, bottom part of the figure). In the following, we describe each of these components.
3.1. Coarse-to-Fine Denoising
As shown by previous works (Hertz et al., 2022; Balaji et al., 2022; Chefer et al., 2023; Patashnik et al., 2023), the denoising process of diffusion models operates in a coarse-to-fine manner: early steps determine low-frequency structures, while high-frequency details emerge in later steps.
This behavior is reflected in the training process, in which varying noise levels are applied across different timesteps, and the model is trained to reconstruct the original clean image from the noisy input. High noise levels used at early timesteps yield low signal-to-noise ratios, restricting reconstruction to spatially smooth, low-frequency components, such as background or layout, which are recoverable even when the image is heavily corrupted. At smaller timesteps, where the signal-to-noise ratio is higher, high-frequency features become more reliably reconstructible.
During inference, the noise is gradually removed as the model generates the image in a coarse-to-fine manner. This coarse-to-fine nature of the image generation leads to two key observations: (i) Irreversibility of details. At each stage of denoising, the model commits to a certain level of structural detail. Once a given attribute (e.g., layout, coarse shape) has been established, it becomes effectively fixed and cannot be revised in later steps, even if it misaligns with the prompt. (ii) Flexibility in high-frequency details. In early stages, high-frequency details have not yet emerged and thus are unaffected by the prompt. This allows greater flexibility in guiding the model without impacting the final generated fine details.
We demonstrate these observations in Figure 3, by examining the model’s $x_{0}$ predictions throughout the denoising process. In the top row, the prompt is fixed to “A howling wolf and a flying bat at midday”. In the middle row, we use the proxy prompt “A howling dog and a flying bird at midday” during the early steps of denoising, and then switch to the same target prompt used in the top row. We replaced “wolf” with “dog” because the two animals are semantically similar and share similar low-frequency visual features. Moreover, “a howling dog” is not contextually associated with nighttime. For similar reasons, we replaced “bat” with “bird”.
The first row illustrates the irreversibility of details. Night colors and a large moon are established early in the denoising process, and although they do not match the prompt, the model is unable to correct them in later steps.
The second column shows the flexibility of high-frequency details. Although different prompts were used for each row, the layout and object shapes are visually similar, reflecting the fact that distinguishing information between the target and proxy prompts – specifically regarding layout and object shape – has not yet emerged at this early stage of the denoising process. Yet, since a howling dog, unlike a howling wolf, is not entangled with night, the background in the second row adheres more closely to the prompt. By switching to the target prompt at this point in the process, we are still able to generate a wolf and a bat rather than a dog and a bird, while also aligning with the word “midday”.
The bottom row illustrates the importance of selecting an appropriate proxy prompt. Snakes and rats are associated with nighttime, which conflicts with the “midday” specification. In addition, their lack of visual similarity to the target objects leads to visual leakage between the bat and the wolf.
These observations motivate a stage-aware prompting strategy that enables precise control over early structural decisions, while allowing more flexible conditioning for attributes that are not yet generated. Building on this, we now describe the two main components of our method: prompt decomposition and stage-aware prompt injection.
Figure 3. Coarse-to-fine denoising. $x_{0}$ predictions at the initial, intermediate, and final time steps. During the early denoising stages, where primarily low-frequency structures are generated, different prompts introducing distinct objects can produce visually similar layouts and objects. Once this structure is established, it becomes effectively frozen and cannot be modified in later steps. As shown in the bottom row, using incorrect placeholder objects in the proxy prompt can anchor erroneous visual information early on, leading to eventual misalignment to the target prompt.
3.2. Prompt Decomposition
Given a prompt $P$ that contains contextually contradictory concepts, we aim to construct a sequence of proxy prompts $\{p_{1},p_{2},...,p_{n}\}$ and corresponding timestep intervals $\{I_{1},I_{2},...,I_{n}\}$ that together reflect the intended semantics of $P$. Each proxy prompt $p_{i}$ is designed to preserve the relevant semantics of $P$ for attributes or concepts that typically form during its assigned interval $I_{i}$. This is crucial due to the irreversibility of details in the generation process. At the same time, $p_{i}$ is chosen to avoid contradictory combinations that tend to emerge at that stage, leveraging the flexibility in high-frequency details. This decomposition enables conditioning the diffusion model on contextually coherent content that evolves in tandem with the coarse-to-fine structure of the denoising process (see Figure 3).
To generate these proxy prompts and their associated intervals, we employ a large language model (LLM) with two complementary roles. First, we leverage the LLM’s broad world knowledge and reasoning capabilities to detect semantic contradictions and to propose contextually appropriate substitutes for conflicting concepts. Second, we use its instruction-following and in-context learning capabilities to infer the temporal staging of concepts, guided by examples and instructions that reflect the denoising behavior of diffusion models.
We implement this using a structured prompt template that includes both instructional text and representative input-output examples, along with textual explanations of the contextual contradictions in the original prompts. These examples illustrate successful decompositions that resolve contextual conflicts, as well as cases that require no decomposition, enabling the LLM to generalize from both scenarios. The full instruction prompt is provided in the appendix (see Table 4).
In-context Examples
Our in-context examples for guiding the LLM include the target prompt as input and the corresponding proxy prompts, each paired with its associated timestep interval, as output. Each example also includes a brief explanation describing the contextual contradiction in the prompt, which is part of the expected output. By requiring the LLM to produce this explanation, we encourage it to engage in a reasoning process, improving its ability to identify contradictions and generate coherent proxy prompts. The examples cover a range of contradiction types, including semantic mismatches, spatial violations, behavioral implausibilities, and stylistic inconsistencies. We provide these examples in the appendix (see Table 5).
These examples were constructed by first identifying contextually-contradictory prompts that produce implausible generations in the base model (FLUX), particularly when concept conflicts appear early in the denoising process. Drawing on our coarse-to-fine analysis, we manually decomposed such prompts into proxy prompts sequences that stabilize low-frequency structure before gradually introducing high-frequency details. We used an LLM to generate natural language explanations and validate that each substitution preserve visual and contextual plausibility while resolving conflicts in a temporally coherent manner.
In total, we include 20 such examples, covering both effective decompositions and cases where minimal intervention was sufficient to preserve coherence. These examples guide the LLM in learning successful decomposition strategies, such as proxy substitution, temporal staging, and contradiction deferral.
Concept Substitution
A core mechanism of our prompt decomposition is concept substitution, which promotes contextual coherence by temporarily replacing a conflicting concept in the prompt with a structurally appropriate alternative (see Figure 3). A simpler alternative for avoiding contextual contradictions is to remove the conflicting concept entirely. However, we observe that introducing an object only at a later stage, without a placeholder proxy, can affect the perceived size of that object (see Figure 4). In extreme cases, introducing the object too late may result in it being omitted from the generated image altogether. Therefore, we prefer the concept substitution strategy.
As demonstrated in Figure 3, choosing the right substitution is key to the method’s success. When the substituted concept is semantically appropriate and visually similar to the original object during the interval in which it is applied, the model correctly associates it with the original concept. In contrast, poor substitutions – due to mismatches in scale, pose, location, or to inappropriate timing – can result in extra objects, layout errors, or failure to follow the prompt (see Figure 3, bottom row, and Figure 5).
“A monkey is sitting on a white sandy beach.”
Refer to caption Refer to caption Refer to caption
“A snowman in a desert.”
Refer to caption Refer to caption Refer to caption
Step 1 Step 3 Step 5
Figure 4. The effect of the object removal strategy. Each image is generated from a prompt that initially describes only the background (“a white sandy beach,” “a desert”). At the timestep indicated below each column, the foreground object (“a monkey,” “a snowman”) is added to the prompt. Removing the object from the initial prompt, without using a placeholder, significantly affects its apparent scale. Early insertion allows the model to allocate more spatial area, while late insertion confines the object to limited regions within the established layout.
3.3. Stage-aware Prompt Injection
Given a sequence of proxy prompts ${p_{1},p_{2},...,p_{n}}$, their corresponding timestep intervals ${I_{1},I_{2},...,I_{n}}$, and a text-to-image (T2I) diffusion model, we condition the model using different prompts throughout the denoising process. At each timestep $t$, we apply the prompt $p_{i}$ such that $t\in I_{i}$. By aligning each proxy prompt with a specific timestep interval, the denoising process is guided at each stage by concepts relevant to the level of detail emerging at that point. This enables the gradual construction of the image while avoiding conflicts with the model’s learned priors.
This injection mechanism integrates seamlessly into existing inference pipelines without any architectural modifications, and is compatible with a range of pretrained diffusion models. Differences in denoising schedules across models are handled entirely during prompt decomposition (Section 3.2). We show examples with SD3 in the appendix (see Figure 11).
“A camping tent is inside a bed room”
Refer to caption Refer to caption Refer to caption
“A mother duck guards three rubber duckies.”
Refer to caption Refer to caption Refer to caption
too early too late Ours
Figure 5. Impact of proxy object timing on generated image attributes. Changing the target object too early fails to disentangle the desired attribute, while modifying it too late prevents effective binding. Attribute control is most successful when applied within the correct time interval. In the top row, the initial proxy prompt is “A pillow fort in a bedroom”, while in the bottom row, the proxy prompt is “A mother duck guards three ducklings”.
4. Experiments and Results
In this section, we conduct qualitative (Section 4.1) and quantitative (Section 4.2) comparisons to assess the effectiveness of our method, along with ablation study (Section 4.3) to evaluate the contribution of individual components. We also present some limitations of our method (Section 4.4).
Implementation Details.
We use the state-of-the-art text-to-image model FLUX.1 [dev] (Labs, 2024) and GPT-4o as LLM to demonstrate the effectiveness of our approach. Inference is performed using $T=50$ sampling steps. In all experiments, we guide the LLM to limit the number of proxy prompts generated during decomposition to a maximum of three per input prompt. We use the baselines hyperparameters as appeared in their original papers or implementation. To ensure a fair comparison, we also use $T=50$ sampling steps for all baseline methods.
Baselines
Since our method builds upon FLUX.1 [dev], we directly compare our results against this base model to clearly demonstrate the improvements achieved by our approach. Additionally, we include a comparison to the training-free method R2F (Park et al., 2024), which similarly employs a large language model (LLM) to alternate between different prompts during generation. As R2F is implemented based on SD3 (Esser et al., 2024), we also compare our results to this model. Lastly, we compare against the fine-tuned model Ella (Hu et al., 2024), which is currently publicly available with SD 1.5 as its base model.
Datasets
We evaluate our method using three datasets: Whoops! (Bitton-Guetta et al., 2023), Whoops-Hard, and ContraBench. Whoops! consists of 500 prompts paired with commonsense-defying images, designed to test visual reasoning and compositional understanding. While relevant to our task, many of its prompts are relatively easy for modern T2I models and fail to consistently reveal their limitations. To address this, we curate a more challenging dataset Whoops-Hard, a subset comprising 100 particularly difficult prompts from Whoops! (Bitton-Guetta et al., 2023). This challenging benchmark provides a stronger basis for evaluating text-to-image generation given the current performance of state-of-the-art methods. Furthermore, to test more challenging cases of semantic alignment, we introduce a new benchmark ContraBench, a curated set of 40 prompts crafted to induce strong contextual contradictions, including background-object mismatches, semantic tension, and scale conflicts. ContraBench serves as a focused, high-difficulty benchmark for evaluating text-image alignment on prompts that contain contextual contradictions. We will make both ContraBench and the Whoops-Hard subset publicly available.
“A monkey juggles tiny elephants.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“Shrek is blue.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“Lightning striking a dilapidated shack on a clear sunny day.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ella SD3 R2F FLUX Ours
Figure 6. Our method effectively resolves contextual contradictions across diverse forms, including inconsistent backgrounds, unnatural color assignments, and implausible actions. Compared to prior approaches, our results exhibit higher semantic coherence and more accurate prompt grounding under contradictory prompts.
4.1. Qualitative Results
Comparisons with baselines.
Figures 6 and 9 present qualitative comparisons on the Whoops! and ContraBench datasets. Across both benchmarks, the baseline methods FLUX, SD3, R2F, and Ella consistently exhibit characteristic failure modes when handling prompts with contextually contradictory concepts. For instance, as shown in Figure 6, our method is the only one to successfully generate scenarios such as a blue Shrek or a monkey juggling tiny elephants.
For SD3 and FLUX, these prompts conflict with their learned priors, leading to failures in generating text-aligned images. Ella was originally designed for long and complex prompts, not specifically for those containing contextual contradictions, and therefore struggles with challenging contradiction cases. The training-free baseline R2F alternates between prompts at predefined timesteps, a strategy originally designed for attribute binding rather than addressing contextual contradictions. While it can reinforce rare concepts, this approach does not align prompt content with the stages at which semantic features emerge during denoising. As a result, it often produces hybrid concepts that merge incompatible elements from conflicting objects (see the owl and SpongeBob in Figure 9).
In contrast, our method produces semantically coherent outputs by introducing proxy prompts at denoising stages where the corresponding visual features are expected to emerge. This alignment enables the model to effectively handle conflicting concepts. Across both Whoops! and ContraBench, our approach consistently generates contextually contradictory elements with high accuracy.
To demonstrate the robustness of our approach, we provide a comparison across several random seeds in Figure 10. The same seeds are used for all methods, showing that our method consistently outperforms the baselines.
Improve realism
Figure 7 demonstrates how our method generates photorealistic and semantically coherent images for prompts with atypical attribute combinations. In contrast, FLUX often defaults to cartoon-like renderings, likely due to priors associating fantastical elements with imaginative visual styles. By using contradiction-free proxy prompts, our method enables the model to produce realistic outputs without explicitly requiring photorealism in the prompt.
“A kangaroo delivering mail on a bicycle.” “A boy is kissing a hedgehog.”
Refer to caption Refer to caption Refer to caption Refer to caption
FLUX
Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Figure 7. FLUX tends to generate realistic images by default. However, when given unrealistic prompts, it often produces cartoon-like samples. In contrast, our method, which gradually resolves such prompts through coherent proxy stages, consistently generates realistic and semantically aligned images.
4.2. Quantitative Results
Our method focuses on solving alignment with challenging and contradictory textual prompts. Therefore, we evaluate its performance by measuring prompt alignment. To do so, we use the GPT-4o vision-language model (VLM). The model is prompted to assign a score from 1 to 5 for each generated image, based on how well the image aligns with the given textual prompt. The exact evaluation prompt is provided in the appendix (See Table 6). For each prompt, we generate outputs using three fixed random seeds, which are applied consistently across all evaluated methods. The average score across the three seeds is then scaled to a range between 20 and 100 by multiplying by 20.
As can be seen in Table 1, our method outperforms the baselines across all three benchmarks. Notably, using SD3 as the base model yields better prompt alignment than the state-of-the-art FLUX model under contradictory prompt settings, while FLUX demonstrates superior visual quality (See Figure 9). In contrast, our method, built on top of FLUX, achieves the highest prompt adherence while preserving its strong visual quality. This results in the best overall performance, combining high-fidelity generation with robust alignment to challenging prompts.
Table 1. Quantitative evaluation on the Whoops dataset using the GPT-4o vision-language model. We report average alignment, where alignment reflects how well the image matches the prompt semantics, without looking at the quality of the image
Benchmarks
Models Whoops Whoops-Hard ContraBench
Ella 69.09 45.19 55.16
SD3.0 82.63 55.73 57.5
R2F (SD3.0) 83.50 57.06 59.16
FLUX 78.85 44.3 57.16
Ours (FLUX) 85.10 62.13 66.16
Table 2. User study results. Win rates of our method in text-image alignment and image quality, compared against each baseline method.
Ella SD3.0 R2F (SD3.0) FLUX
Alignment $81\%$ $70\%$ $75\%$ $81\%$
Quality $74\%$ $72\%$ $68\%$ $63\%$
User study
VLM-based automatic metrics, such as those shown in Table 1, often fail to detect subtle semantic inconsistencies, particularly when evaluating generations conditioned on complex or contradictory prompts. Moreover, these models do not adequately assess the visual quality of the generated images. To address these limitations, we conducted a user study aimed at evaluating both prompt adherence and visual quality from a human perspective. We randomly sampled 24 prompts from the Whoops! and ContraBench benchmarks. For each prompt, participants were shown two images: one generated by our method and the other by one of the baselines. They were asked to answer two questions: (1) which image most accurately reflected the prompt, and (2) which image had better visual quality.
In total, we collected responses from 61 users, resulting in 1,464 individual evaluations. The results of the user study are summarized in Table 2, which reports the win rate of our method when compared pairwise against each baseline. These results highlight the superiority of our approach in handling contextually contradictory prompts, achieving both better prompt alignment and visual quality compared to baseline approaches.
Refer to caption Refer to caption Refer to caption
“A bouquet of flowers is upside down in a vase” “A white glove has 6 fingers” “The shadow of a cat is facing the opposite direction”
Figure 8. Failure cases from Whoops!.
4.3. Ablation Studies
To evaluate the contribution of each component in our method, we perform an ablation study on the Whoops-Hard benchmark by selectively removing or modifying individual elements of the method, and present the results in Table 3. Each configuration isolates a specific design choice to quantify its effect on alignment performance. We observe that using in-context examples significantly improves the LLM’s ability to decompose and interpret contextually contradictory prompts, resulting in better text-image alignment. Removing the explanation requirement from the LLM impairs its reasoning capabilities and leads to a notable performance drop, suggesting that requiring the LLM to generate explicit explanations helps it make more coherent semantic decisions, ultimately improving prompt decomposition. Additionally, using only two proxy prompts, underperforms compared to the full method, highlighting the importance of introducing prompt information at targeted stages for robust alignment.
4.4. Limitations
In Figure 8, we present three failure cases of our method on examples from the Whoops! benchmark. Since our approach relies on guiding the model through text alone, it cannot control properties that the underlying model inherently struggles with—such as generating specific quantities or enforcing precise orientations.
Table 3. Ablation study on the Whoops-Hard Benchmark. We evaluate the impact of key components of our prompt decomposition method by: (1) removing in-context examples, (2) removing the explanation requirement, and (3) restricting the decomposition to at most two proxy prompts. Results show a significant drop in alignment when any of these components are removed.
static w/o in-context w/o reasoning 2 proxy Full
Alignment 44.3 48.0 46.46 50.73 62.13
5. Conclusions
We have introduced a training-free framework for resolving contextual contradictions in text-to-image generation – that is, cases where seemingly plausible prompts are internally inconsistent because of strong, hidden biases in the model. Our method identifies and strategically separates such contradictions over time, enabling faithful rendering of prompts that would otherwise yield semantically inconsistent outputs.
At its core, our approach leverages the broad real-world understanding embedded in large language models to reason about the semantic plausibility of prompts. Rather than modifying the training data or fine-tuning the model, we harness this reasoning capacity at inference time to decompose prompts into a sequence of contextually coherent stages. These proxy prompts guide the generative process in alignment with the model’s prior expectations, helping it circumvent conflicts and maintain semantic fidelity.
Extensive experiments demonstrate that our stage-aware decomposition substantially improves alignment and image quality – especially for prompts that challenge the model’s default associations. Future directions include refining our method using VLM-generated responses to improve proxy prompts generation dynamically during the generation process, and applying our method to text-to-video generation models.
Appendix A Appendix
Table 4. Full LLM prompt instruction for time step dependent prompt conditioning, used to decompose prompts by denoising stage.
<System Prompt>
You are an expert assistant in time step dependent prompt conditioning for diffusion models.
Your task is to decompose a complex or contextually contradictory prompt into up to three intermediate prompts that align with the model’s denoising stages — from background layout to object identity to fine detail. Only introduce prompt transitions when needed.
Diffusion Semantics (Low $\rightarrow$ High Frequency Progression):
Steps 0–2: Scene layout and dominant color regions (e.g., sky, forest, sand tone)
Steps 3–6: Object shape, size, pose, and position
Steps 7–10: Object identity, material, and surface type (e.g., glass vs. rubber)
Steps 11–13+: Fine features and local details (e.g., tattoos, insects, facial detail)
Since denoising progresses from coarse to fine, it is crucial to stabilize large-scale visual structures (such as body shape, pose, and background) before introducing small or semantically charged elements (such as facial details, objects in hand, or surreal features).
Substitution Strategy:
1. Begin with high-level layout (background, geometry).
2. Use placeholder concepts if needed to stabilize layout before detailed insertions.
3. Substitutes must match in shape, size, and visual function.
4. Replace placeholders as soon as fidelity permits.
5. Do not maintain substitutions longer than needed.
6. If the prompt is visually coherent, return a single prompt with no decomposition.
Output Format:
a. Explanation: A short sentence explaining why decomposition is needed.
b. Final dictionary:
  "prompts_list": [
    "<prompt1>",
    "<prompt2>",
    "<target prompt>"
  "switch_prompts_steps": [<step1>, <step2>, ...]
- The length of switch_ prompts_steps should be one less than prompts_list. - Do not include any text outside this structure.
A.1. LLM Instruction for Prompt Decomposition
Tables 4 and 5 detail the full LLM instruction used for our method’s decomposition, along with the corresponding in-context examples. In a single inference pass, our method detects contextual contradictions, generates proxy prompts, and assigns timestep intervals.
Table 8 presents examples of our LLM input prompts, along with the corresponding output explanations and the decomposition into proxy prompts and timestep intervals.
Table 5. In-context examples for time step dependent prompt conditioning, illustrating reasoning steps and prompt decomposition across denoising stages.
Input: A polar bear in a desert
Output:
a. Explanation: A polar bear is common in snowy scenes, not deserts. Since no suitable object proxy exists, the prompt starts with the desert alone before introducing the unlikely animal.
b. Final dictionary:
  "prompts_list": [
    "A desert",
    "A polar bear in a desert"
  "switch_prompts_steps": [2]
Input: A fox in a nursery
Output:
a. Explanation: A fox is uncommon in indoor scenes. Starting with a dog, then a visually similar breed (Shiba Inu), provides a natural proxy before introducing the fox in a childlike setting.
b. Final dictionary:
  "prompts_list": [
    "A dog in a nursery",
    "A Shiba Inu dog in a nursery",
    "A fox in a baby room"
  "switch_prompts_steps": [4, 7]
Input: A shoes tree in the meadow
Output:
a. Explanation: Treating a “shoes tree” like an “apple tree” allows the model to build structure and object layout using familiar patterns. Introducing apples first grounds the scene in a biologically plausible layout before shifting to the surreal concept of shoes as fruit-like objects.
b. Final dictionary:
  "prompts_list": [
    "tree in the meadow",
    A tree full of apples, in the meadow",
    "A tree full of shoes in the meadow"
  "switch_prompts_steps": [4,8]
Input: A red sports car parked on a mountain road
Output:
a. Explanation: This is a realistic and visually coherent scene, so no decomposition is needed.
b. Final dictionary:
  "prompts_list": [
    "A red sports car parked on a mountain road"
  "switch_prompts_steps": []
A.2. VLM evaluation
We evaluate text-to-image alignment for the different models on the benchmarks using the VLM GPT-4o. The instruction prompt used for the VLM is presented in Table 6.
A.3. ContraBench
To evaluate image generation under prompts containing contextual contradictions, we introduce ContraBench, presented in Table 7. This dataset consists of 40 prompts generated using ChatGPT, each intentionally constructed to include visual or semantic inconsistencies challenge standard diffusion models. The prompts span a range of contradiction types, including physical implausibility, role violation, object misplacement, and surreal transformations. ContraBench serves as a targeted benchmark for assessing a model’s ability to produce visually coherent and semantically plausible images in the presence of contradictory concepts.
A.4. Method Generalization to Different Models
Our method’s proxy prompt injection mechanism integrates seamlessly into existing inference pipelines without requiring any architectural modifications, and it is compatible with a range of pretrained diffusion models. Differences in denoising schedules across models are handled entirely during the prompt decomposition stage. See Figure 11 for an example of how our method integrates with SD3 without any modifications.
Table 6. VLM instruction for prompt alignment. Instructions provided to the VLM for scoring semantic alignment of generated images.
You are an assistant evaluating an image on how well it aligns with the meaning of a given text prompt.
The text prompt is: "{prompt}"
PROMPT ALIGNMENT (Semantic Fidelity)
Evaluate only the meaning conveyed by the image — ignore visual artifacts.
Focus on:
  • Are the correct objects present and depicted in a way that clearly demonstrates their intended roles and actions from the prompt?
  • Does the scene illustrate the intended situation or use-case in a concrete and functional way, rather than through symbolic, metaphorical, or hybrid representation?
  • If the described usage or interaction is missing or unclear, alignment should be penalized.
  • Focus strictly on the presence, roles, and relationships of the described elements — not on rendering quality.
Score from 1 to 5:
  • 5: Fully conveys the prompt’s meaning with correct elements
  • 4: Mostly accurate — main elements are correct, with minor conceptual or contextual issues
  • 3: Main subjects are present but important attributes or actions are missing or wrong
  • 2: Some relevant components are present, but key elements or intent are significantly misrepresented
  • 1: Does not reflect the prompt at all
Respond using this format:
### ALIGNMENT SCORE: <score>
### ALIGNMENT EXPLANATION: <explanation>
SD3 Ours(SD3) SD3 Ours(SD3)
Refer to caption Refer to caption Refer to caption Refer to caption
“A grown man has a baby’s pacifier in his mouth.” “A camping tent is inside a bed room.”
Refer to caption Refer to caption Refer to caption Refer to caption
“A mouse eats a snake.” “A dragon is blowing water.”
Figure 11. Generated images from our method applied on SD3.
Table 7. ContraBench prompt set.
ID Prompt
1 A professional boxer does a split
2 A bear performing a handstand in the park
3 A photorealistic photo of SpongeBob SquarePants dancing ballet
4 A cowboy swimming competitively in an Olympic pool
5 A cruise ship parked in a bathtub
6 A man giving a piggyback ride to an elephant
7 A zebra climbing a tree
8 A coffee machine dispensing glitter
9 A vending machine in a human running posture
10 A ballerina aggressively flipping a table
11 A bathtub floating above a desert in a tornado
12 A monkey juggles tiny elephants
13 a woman has a marine haircut
14 A tower with two hands
15 An archer is shooting flowers with a bow
16 A baseball player backswing a yellow ball with a golf club
17 A barn built atop a skyscraper rooftop
18 A cat balancing a skyscraper on its nose
19 A cow grazing on a city rooftop
20 A fireplace burning inside an igloo
21 A mosquito pulling a royal carriage through Times Square
22 A grandma is ice skating on the roof
23 A muscular ferret in the woods
24 A house with a circular door
25 A photorealistic image of a bear ironing clothes in a laundry room
26 A pizza being used as an umbrella in the rain
27 A cubist lion hiding in a photorealistic jungle
28 A chicken is smiling
29 A realistic photo of an elephant wearing slippers
30 A computer mouse eating a piece of cheese
31 A horse taking a selfie with a smartphone
32 A sheep practicing yoga on a mat
33 a snake eating a small golden guitar
34 A soccer field painted on a grain of rice
35 A snake with feet
36 A woman brushing her teeth with a paintbrush
37 A horse with a hump
38 A hyperrealistic unicorn made of origami
39 A library printed on a butterfly’s wings
40 A bodybuilder balancing on pointe shoes
Table 8. Examples of prompt decomposition for contradictory prompts using our method.
Original Prompt Explanation Prompt Sequence Switch Step(s)
A grown man has a baby’s pacifier in his mouth. The grown man is established first for structure, while the pacifier, being a small but semantically significant detail, is added once the main identity is resolved. (1) A grown man with a small object in his mouth (2) A grown man has a baby’s pacifier in his mouth 4
A dragon is blowing water. Dragons are more commonly depicted blowing fire. A proxy of white smoke is visually similar to water mist in texture, stabilizing the emission process before resolving the surreal water emission. (1) A dragon blowing white smoke (2) A dragon blowing water 3
A pizza with grape toppings. Pizza with traditional toppings stabilizes the geometry and color before introducing the visually similar yet unusual grape topping. (1) A pizza with pepperoni toppings (2) A pizza with grape toppings 3
A coin floats on the surface of the water. Coins typically sink in water, not float. Starting with a leaf—an object that naturally floats—ensures that this behavior within the scene is handled correctly before introducing the coin. (1) A leaf floats on the surface of the water (2) A coin floats on the surface of the water 4
A cockatoo parrot swimming in the ocean. Cockatoos are birds and naturally do not swim; starting with a simple bird on water stabilizes position and motion. Progressing to a duck, before introducing the cockatoo parrot, eases the transition into the final surreal visual. (1) A duck swimming in the ocean (2) A parrot swimming in the ocean (3) A cockatoo parrot swimming in the ocean 3, 6
A woman writing with a dart. Writing is typically associated with a pen or pencil. Starting with a pen stabilizes the action before introducing the unusual tool. (1) A woman writing with a pen (2) A woman writing with a dart 3
Shrek is blue. Shrek is a distinct character with a recognizable green color. Using a simple b̈lue ogreïnitially sets the stage for a color change before fully introducing Shrek to ensure visual coherence. (1) A blue ogre (2) Shrek is blue 3

### v2

@@
-Image Generation from Contextually-Contradictory Prompts
-Abstract.
+Image Generation from Contextually Contradictory Prompts
+Abstract
 Text-to-image diffusion models excel at generating high-quality, diverse images from natural language prompts.
@@
 Experiments across a variety of challenging prompts show substantial improvements in alignment to the textual prompt.
-copyright: nonedoi: XXXXXXX.XXXXXXXsubmissionid: 1201
-Figure 1.
-Our method addresses contextual contradictions in text-to-image generation.
+Figure 1: Our method addresses contextual contradictions in text-to-image generation.
 These contradictions arise when one concept implicitly conflicts with another due to the model’s learned associations.
-For example, if a concept like “nest” is strongly entangled with “birds”, it may conflict with another concept in the prompt such as “rooster”, leading the model to ignore or distort part of the semantic meaning.
-1. Introduction
-Text-to-image diffusion models have demonstrated remarkable capabilities in generating high-quality and diverse visual content from natural language prompts (Rombach et al., 2022; Ramesh et al., 2021; Ho et al., 2020).
+For example, if a concept like “butterfly” is strongly entangled with “flowers”, it may conflict with another concept in the prompt such as “bee’s hive”, leading the model to ignore or distort part of the semantic meaning.
+1 Introduction
+Text-to-image diffusion models have demonstrated remarkable capabilities in generating high-quality and diverse visual content from natural language prompts [27, 24, 15].
 However, achieving precise semantic alignment between the generated image and the conditioning prompt remains an open challenge.
 This issue becomes particularly pronounced when the target prompt lies outside the model’s training distribution, such as prompts that combine semantically plausible yet statistically uncommon or unconventional concepts.
-For example, as shown in Figure 1, generating an image from the prompt “a rooster in a nest” often results in colorful birds sitting in a nest.
-This is due to the model’s strong prior entanglement between bird species and nests, which implicitly contradicts the notion of a rooster.
+For example, as shown in Figure 1, generating an image from the prompt “A butterfly is in a bee’s hive” often results in a butterfly on a flower.
+This is due to the model’s prior that entangles butterflies with flowers, which implicitly contradicts the notion of a bee’s hive.
 We refer to this phenomenon as Contextual Contradiction, a conflict between two concepts that arises not from direct semantic opposition, but from the model’s associations learned during training.
-More precisely, we say that concept $A$ contextually contradicts concept $B$ if the model’s prior entangles $A$ with concept $C$, and $B$ contradicts $C$.
+More precisely, we say that concept $A$ contextually contradicts concept $B$ if the model’s prior entangles $A$ with concept $C$, and $B$ contradicts $C$ (see Figure 2).
 In Figure 1, we illustrate this with a blowing dragon, which stands in contextual contradiction with the water due to its entanglement with fire.
+Figure 2: On the left is a direct contradiction, since a girl cannot smile and frown at the same time.
+On the right is a contextual contradiction: while Disneyland and crying are not directly opposed, the model’s prior associates Disneyland with happiness, which conflicts with crying.
 The phenomenon of contextual contradiction in text-to-image models relates to the broader issue of spurious correlations in deep learning.
-Models often rely on shortcuts, which are predictive patterns in the training data that do not reflect true semantics (Geirhos et al., 2020).
-In NLP, models exploit syntactic heuristics rather than reasoning (McCoy et al., 2019), while in computer vision, they may depend on background cues rather than object features (Xiao et al., 2021; Singh et al., 2020).
-For example, classifiers may fail to recognize cows on beaches due to background biases (Beery et al., 2018).
-In this paper, we identify similar biases in generative models, where contextual contradictions arise when prompts include plausible but uncommon concept combinations.
-To address this issue, we build on the observation that the denoising process follows a coarse-to-fine progression, during which different semantic attributes (e.g., background, pose, shape, and texture) emerge at distinct stages (Chefer et al., 2023; Balaji et al., 2022; Patashnik et al., 2023).
+Models often exploit shortcuts, relying on correlations in the training data that are statistically strong but semantically misleading [13].
+In this paper, we identify a similar bias in generative models: contextual contradictions occur when prompts combine concepts that individually align with the model’s priors but conflict when combined, revealing the model’s reliance on such correlations rather than robust compositional reasoning.
+To address this issue, we introduce Stage-Aware Prompting (SAP), which builds on the observation that the denoising process follows a coarse-to-fine progression, during which different semantic attributes (e.g., background, pose, shape, and texture) emerge at distinct stages [9, 4, 23].
 Our key idea is to guide the model at each stage of denoising with the information most relevant to the type of content being formed at that point.
@@
 For example, it involves understanding that a bear is entangled with specific poses, such as walking on all fours or standing upright, which in turn, contradicts the handstand pose.
-Our method leverages a large language model (LLM) to analyze the target prompt, identify contextual contradictions, and construct suitable proxy prompts.
+To achieve this, SAP leverages a large language model (LLM) to analyze the target prompt, identify contextual contradictions, and construct suitable proxy prompts.
 We demonstrate that, by using in-context examples and prompting the LLM to follow a reasoning process through a brief explanation, it can identify contextually contradictory concepts in a prompt and determine the appropriate stage of denoising at which each attribute should be introduced.
@@
 In doing so, the LLM effectively guides the model toward the intended meaning of the original prompt.
-Through extensive experiments, we demonstrate the effectiveness of our method in generating images from contextually contradicting text prompts.
-By introducing prompt information at targeted stages, our approach generates precise combination of semantic attributes while avoiding undesired entanglement of unrelated attributes.
-Compared to previous methods, our stage-dependent prompt decomposition leads to more faithful and semantically aligned generations.
-2. Related Work
-Learned Spurious Correlations
-Machine learning models are known to be sensitive to spurious correlations in their training data (Geirhos et al., 2020; McCoy et al., 2019; Ye et al., 2024).
-This sensitivity often leads to a significant drop in performance when training-time associations between concepts do not hold at test time.
-Prior work has extensively studied this issue in discriminative computer vision tasks, showing, for example, that recognition models tend to rely heavily on background cues (Singh et al., 2020; Xiao et al., 2021; Beery et al., 2018).
+Through extensive experiments, we demonstrate the effectiveness of SAP in generating images from contextually contradictory text prompts.
+By introducing prompt information at targeted stages, SAP generates precise combinations of semantic attributes while avoiding undesired entanglement.
+Compared to previous methods, SAP’s stage-dependent prompt decomposition leads to more faithful and semantically aligned generations.
+2 Related Work
+Learned Spurious Correlations.
+Machine learning models are known to be sensitive to spurious correlations in their training data [13, 20, 33], leading to performance drops when training-time associations do not hold at test time.
+Prior work has extensively studied this in discriminative vision tasks, showing, for example, that recognition models tend to rely heavily on background cues [28, 31, 6].
 In our work, we show that text-to-image models exhibit similar behavior.
@@
 We evaluate this behavior using the Whoops!
-dataset (Bitton-Guetta et al., 2023), which contains prompts constructed by first describing two co-occurring elements, and then replacing one with a less compatible alternative.
+dataset [7], which contains prompts constructed by first describing two co-occurring elements, and then replacing one with a less compatible alternative.
 This results in scenarios that are unlikely to occur in the real world.
-Semantic Alignment in Text-to-Image Synthesis
+Semantic Alignment in Text-to-Image Synthesis.
 Text-to-image models often struggle to fully capture the semantic intent of input prompts, particularly when prompts involve complex or internally conflicting concepts.
-Previous works have analyzed common failure cases and proposed targeted improvements across various stages of the generation pipeline, including enhanced text embedding representations (Rassin et al., 2022; Feng et al., 2022; Tunanyan et al., 2023), refined attention mechanisms (Dahary et al., 2024) and guidance strategies that leverage attention maps for loss heuristics (Chefer et al., 2023; Rassin et al., 2023; Meral et al., 2024; Agarwal et al., 2023).
+Previous works have analyzed common failure cases and proposed targeted improvements across various stages of the generation pipeline, including enhanced text embedding representations [26, 12, 29], refined attention mechanisms [10], guidance strategies that leverage attention maps for loss heuristics [9, 25, 21, 2] and dynamic guidance scheduling via annealed classifier-free guidance [34].
 Despite these advances, existing methods often fail to handle prompts containing contradictory concepts arising from the model’s learned associations.
 Our work directly addresses this underexplored challenge, focusing on contextually contradictory prompts.
-Multi-Prompt Conditioning Techniques
+“A howling wolf.” “A polar bear.” “A dragon blowing.”
+Refer to caption Refer to caption Refer to caption
+<howling> <polar> <dragon>
+Refer to caption Refer to caption Refer to caption
+(a) Entangled concepts.
+“A woman writing.” “A woman writing with a dart.”
+Refer to caption Refer to caption Refer to caption
+<dart> <writing> log ratio
+Refer to caption Refer to caption Refer to caption
+(b) From entanglement to contradiction.
+Figure 3: (a) By examining attention maps, we observe that textual tokens embed contextual associations, leading to the generation of concepts not explicitly mentioned in the prompt.
+For example, the token ‘howling’ encourages the presence of the moon, as indicated by its strong attention connection.
+(b) In prompts with contextual contradictions, two tokens may overlap in attending to the same region, as seen with ‘dart’ and ‘writing’.
+The token ‘writing’ dominates this region, as shown in the log-ratio map, where red areas indicate stronger attention to ‘writing’ relative to ‘dart’.
+Multi-Prompt Conditioning Techniques.
 Conditioning diffusion models on multiple prompts has emerged as an effective strategy for improving control and compositionality.
-One line of work, primarily focused on personalization, introduces distinct learned tokens at different layers of the model and at various denoising timesteps (Alaluf et al., 2023; Voynov et al., 2023).
+One line of work, primarily focused on personalization, introduces distinct learned tokens at different layers of the model and at various denoising timesteps [3, 30].
 This design allows each token to capture different attributes of the personalized concept, leading to improved identity preservation.
-Other approaches vary the prompt across timesteps to modulate specific visual properties, such as object shape (Liew et al., 2022; Patashnik et al., 2023), or alternate between rare and frequent object descriptions to improve attribute binding (Park et al., 2024).
-Fine-grained spatial control has also been achieved by assigning sub-prompts to separate image regions (Yang et al., 2024).
-Additionally, some methods leverage multiple diffusion models, each conditioned on different prompt attributes, and combine their outputs into a unified prediction (Liu et al., 2022; Bar-Tal et al., 2023).
-Unlike most prior works, we focus on utilizing multiple prompt to settle internal semantic tension, where concept combinations lead to contextual contradictions.
-Figure 2.
-Our method navigates the denoising process using time-dependent proxy prompts.
-(Top) A large language model (LLM) is employed to decompose a target prompt with entangled or conflicting attributes into a sequence of proxy prompts and their corresponding time-step intervals.
-(Bottom) These intermediate prompts are injected into the diffusion process at the specified intervals.
-This time-dependent prompt conditioning aligns more closely with the model’s prior at each stage, enabling a semantically consistent and temporally coherent progression through the denoising process.
-[…]
-LLM-Guided Diffusion
+Other approaches vary the prompt across timesteps to modulate specific visual properties, such as object shape [18, 23], or alternate between rare and frequent object descriptions to improve attribute binding [22].
+Fine-grained spatial control has also been achieved by assigning sub-prompts to separate image regions [32].
+Additionally, some methods leverage multiple diffusion models, each conditioned on different prompt attributes, and combine their outputs into a unified prediction [19, 5].
+Unlike most prior works, we focus on utilizing multiple prompts to settle internal semantic tension, where concept combinations lead to contextual contradictions.
+LLM-Guided Diffusion.
 LLMs have demonstrated strong capabilities in language understanding.
-Moreover, they possess a broad, implicit grasp of world knowledge, acquired through large-scale training on diverse textual data.
-Recent approaches have leveraged these capabilities to guide diffusion model generation, often incorporating planning and reasoning to improve semantic alignment (Yang et al., 2024; Park et al., 2024; Hu et al., 2024).
-In our work, we employ LLMs with in-context learning to identify contextual contradictions, generate proxy prompts, and determine the corresponding timestep ranges for conditioning, while promoting a reasoning process through brief, explanatory outputs.
-3. Method
+They also capture broad world knowledge through large-scale training on diverse text.
+Recent approaches have leveraged these capabilities to guide diffusion model generation, often incorporating planning and reasoning to improve semantic alignment [32, 22, 16].
+In our work, we employ LLMs with in-context learning to identify contextual contradictions, generate proxy prompts, and determine the corresponding timestep ranges for conditioning, while encouraging reasoning through brief explanatory outputs.
+3 From Entangled Concepts to Contextual Contradictions
+Diffusion models inherit strong distributional biases from their training data, where objects are frequently tied to specific contexts.
+For example, prompts like “a duck” almost always result in water backgrounds, and “a polar bear” appears in snow.
+These reflect entangled concepts, learned associations that go beyond the explicit text.
+While often helpful, such priors hinder generation when prompts require unusual or contradictory combinations.
+To study this effect, we analyze attention maps (Figure 3(a)), which indicate the spatial regions influenced (attended) by each token.
+We find that text tokens often attend not only to the image regions directly corresponding to the object but also to contextually linked elements.
+For example, in “a howling wolf”, the token ‘howling’ influences both the mouth and the moon.
+Similarly, ‘dragon’ attends to flames even when fire is not mentioned.
+These patterns reveal that the model encodes distributional correlations beyond literal semantics.
+These distributional correlations contribute to the difficulty of generating images with contextual contradictions.
+For example, when generating an image from the prompt “a woman writing with a dart”, the model fails to replace the pen with a dart.
+The attention maps shed light on this failure: both ‘writing’ and ‘dart’ attend to the same spatial area (the hand/tool), but ‘writing’ dominates (see log-ratio maps in Figure 3(b)), suppressing the influence of ‘dart’.
+This reflects a failure mode in which entrenched associations override less familiar ones, preventing proper integration of conflicting concepts.
+4 Stage Aware Prompting (SAP)
+Figure 4: Coarse-to-fine denoising with stage-aware prompting.
+We show $x_{0}$ predictions at initial, intermediate, and final steps.
+A is the target prompt, B a suitable proxy, and C an unsuitable proxy.
+Using A alone locks in night and moon despite “midday”.
+Using B first and then switching to A preserves a daytime layout and later adapts the identities to wolf and bat.
+Using C first sets a layout without a flying object, so the final image fails to produce the intended subjects.
 In this section, we begin by analyzing the coarse-to-fine behavior of the denoising process (Section 4.1).
@@
 In the following, we describe each of these components.
-3.1. Coarse-to-Fine Denoising
-As shown by previous works (Hertz et al., 2022; Balaji et al., 2022; Chefer et al., 2023; Patashnik et al., 2023), the denoising process of diffusion models operates in a coarse-to-fine manner: early steps determine low-frequency structures, while high-frequency details emerge in later steps.
-This behavior is reflected in the training process, in which varying noise levels are applied across different timesteps, and the model is trained to reconstruct the original clean image from the noisy input.
-High noise levels used at early timesteps yield low signal-to-noise ratios, restricting reconstruction to spatially smooth, low-frequency components, such as background or layout, which are recoverable even when the image is heavily corrupted.
-At smaller timesteps, where the signal-to-noise ratio is higher, high-frequency features become more reliably reconstructible.
-During inference, the noise is gradually removed as the model generates the image in a coarse-to-fine manner.
-This coarse-to-fine nature of the image generation leads to two key observations: (i) Irreversibility of details.
-At each stage of denoising, the model commits to a certain level of structural detail.
-Once a given attribute (e.g., layout, coarse shape) has been established, it becomes effectively fixed and cannot be revised in later steps, even if it misaligns with the prompt.
+4.1 Coarse-to-Fine Denoising
+Diffusion models generate images in a coarse-to-fine manner: early steps determine low-frequency structures, while high-frequency details emerge in later steps [14, 4, 9, 23].
+From this behavior, we draw two key observations: (i) Irreversibility of details.
+As denoising progresses, the model sequentially commits to different levels of detail, beginning with layout and overall shape.
+Once these are formed, they cannot be revised in later stages, even if they conflict with the prompt.
 (ii) Flexibility in high-frequency details.
-In early stages, high-frequency details have not yet emerged and thus are unaffected by the prompt.
-This allows greater flexibility in guiding the model without impacting the final generated fine details.
-We demonstrate these observations in Figure 3, by examining the model’s $x_{0}$ predictions throughout the denoising process.
-In the top row, the prompt is fixed to “A howling wolf and a flying bat at midday”.
-In the middle row, we use the proxy prompt “A howling dog and a flying bird at midday” during the early steps of denoising, and then switch to the same target prompt used in the top row.
-We replaced “wolf” with “dog” because the two animals are semantically similar and share similar low-frequency visual features.
-Moreover, “a howling dog” is not contextually associated with nighttime.
-For similar reasons, we replaced “bat” with “bird”.
-The first row illustrates the irreversibility of details.
-Night colors and a large moon are established early in the denoising process, and although they do not match the prompt, the model is unable to correct them in later steps.
-The second column shows the flexibility of high-frequency details.
-Although different prompts were used for each row, the layout and object shapes are visually similar, reflecting the fact that distinguishing information between the target and proxy prompts – specifically regarding layout and object shape – has not yet emerged at this early stage of the denoising process.
-Yet, since a howling dog, unlike a howling wolf, is not entangled with night, the background in the second row adheres more closely to the prompt.
-By switching to the target prompt at this point in the process, we are still able to generate a wolf and a bat rather than a dog and a bird, while also aligning with the word “midday”.
-The bottom row illustrates the importance of selecting an appropriate proxy prompt.
-Snakes and rats are associated with nighttime, which conflicts with the “midday” specification.
-In addition, their lack of visual similarity to the target objects leads to visual leakage between the bat and the wolf.
-These observations motivate a stage-aware prompting strategy that enables precise control over early structural decisions, while allowing more flexible conditioning for attributes that are not yet generated.
+In early stages, high-frequency details are absent and unaffected by the prompt, enabling flexible guidance without influence on fine details.
+As shown in Section 3, contextual contradictions stem from concept entanglement in the diffusion prior.
+Since different concepts emerge at different levels of detail during the coarse-to-fine denoising process, they can be decomposed across the denoising stages.
+We illustrate this in Figure 4 by examining the model’s $x_{0}$ predictions across denoising steps.
+In the top-left row, the prompt “a howling wolf and a flying bat at midday” shows that early steps already impose entangled nighttime and moon structures, contradicting “midday”.
+In the bottom-left row, starting with a proxy prompt containing “dog” and “bird” (instead of “wolf” and “bat”, respectively) and later switching to the target prompt produces a correct midday scene with the intended objects.
+The bottom-left row demonstrates both the irreversibility of coarse details (the scene remains daytime) and the flexibility of high-frequency details (the object identities adapt).
+The top-right row highlights the importance of selecting an appropriate proxy prompt.
+Since rats do not fly, the layout determined in early steps lacks a flying object, resulting in a sitting wolf with bat wings.
+These observations motivate a stage-aware prompting strategy that fixes structural decisions early while keeping later attributes flexible.
 Building on this, we now describe the two main components of our method: prompt decomposition and stage-aware prompt injection.
-Figure 3.
-Coarse-to-fine denoising.
-$x_{0}$ predictions at the initial, intermediate, and final time steps.
-During the early denoising stages, where primarily low-frequency structures are generated, different prompts introducing distinct objects can produce visually similar layouts and objects.
-Once this structure is established, it becomes effectively frozen and cannot be modified in later steps.
-As shown in the bottom row, using incorrect placeholder objects in the proxy prompt can anchor erroneous visual information early on, leading to eventual misalignment to the target prompt.
-3.2. Prompt Decomposition
+4.2 Prompt Decomposition
+Figure 5: SAP generates images from contextually contradictory prompts using time-dependent proxy prompts.
+(Top) A large language model (LLM) decomposes the target prompt into a sequence of proxy prompts with corresponding timestep intervals.
+(Bottom) These proxy prompts are injected into the diffusion process at their designated intervals to guide generation.
 Given a prompt $P$ that contains contextually contradictory concepts, we aim to construct a sequence of proxy prompts $\{p_{1},p_{2},...,p_{n}\}$ and corresponding timestep intervals $\{I_{1},I_{2},...,I_{n}\}$ that together reflect the intended semantics of $P$.
-Each proxy prompt $p_{i}$ is designed to preserve the relevant semantics of $P$ for attributes or concepts that typically form during its assigned interval $I_{i}$.
-This is crucial due to the irreversibility of details in the generation process.
-At the same time, $p_{i}$ is chosen to avoid contradictory combinations that tend to emerge at that stage, leveraging the flexibility in high-frequency details.
-This decomposition enables conditioning the diffusion model on contextually coherent content that evolves in tandem with the coarse-to-fine structure of the denoising process (see Figure 3).
-To generate these proxy prompts and their associated intervals, we employ a large language model (LLM) with two complementary roles.
-First, we leverage the LLM’s broad world knowledge and reasoning capabilities to detect semantic contradictions and to propose contextually appropriate substitutes for conflicting concepts.
-Second, we use its instruction-following and in-context learning capabilities to infer the temporal staging of concepts, guided by examples and instructions that reflect the denoising behavior of diffusion models.
-We implement this using a structured prompt template that includes both instructional text and representative input-output examples, along with textual explanations of the contextual contradictions in the original prompts.
-These examples illustrate successful decompositions that resolve contextual conflicts, as well as cases that require no decomposition, enabling the LLM to generalize from both scenarios.
-The full instruction prompt is provided in the appendix (see Table 4).
-In-context Examples
-Our in-context examples for guiding the LLM include the target prompt as input and the corresponding proxy prompts, each paired with its associated timestep interval, as output.
-Each example also includes a brief explanation describing the contextual contradiction in the prompt, which is part of the expected output.
-By requiring the LLM to produce this explanation, we encourage it to engage in a reasoning process, improving its ability to identify contradictions and generate coherent proxy prompts.
-The examples cover a range of contradiction types, including semantic mismatches, spatial violations, behavioral implausibilities, and stylistic inconsistencies.
-We provide these examples in the appendix (see Table 5).
-These examples were constructed by first identifying contextually-contradictory prompts that produce implausible generations in the base model (FLUX), particularly when concept conflicts appear early in the denoising process.
-Drawing on our coarse-to-fine analysis, we manually decomposed such prompts into proxy prompts sequences that stabilize low-frequency structure before gradually introducing high-frequency details.
-We used an LLM to generate natural language explanations and validate that each substitution preserve visual and contextual plausibility while resolving conflicts in a temporally coherent manner.
-In total, we include 20 such examples, covering both effective decompositions and cases where minimal intervention was sufficient to preserve coherence.
-These examples guide the LLM in learning successful decomposition strategies, such as proxy substitution, temporal staging, and contradiction deferral.
-Concept Substitution
-A core mechanism of our prompt decomposition is concept substitution, which promotes contextual coherence by temporarily replacing a conflicting concept in the prompt with a structurally appropriate alternative (see Figure 3).
-A simpler alternative for avoiding contextual contradictions is to remove the conflicting concept entirely.
-However, we observe that introducing an object only at a later stage, without a placeholder proxy, can affect the perceived size of that object (see Figure 4).
-In extreme cases, introducing the object too late may result in it being omitted from the generated image altogether.
-Therefore, we prefer the concept substitution strategy.
-As demonstrated in Figure 3, choosing the right substitution is key to the method’s success.
-When the substituted concept is semantically appropriate and visually similar to the original object during the interval in which it is applied, the model correctly associates it with the original concept.
-In contrast, poor substitutions – due to mismatches in scale, pose, location, or to inappropriate timing – can result in extra objects, layout errors, or failure to follow the prompt (see Figure 3, bottom row, and Figure 5).
+Each proxy prompt $p_{i}$ is designed to (i) preserve the relevant semantics of $P$ for attributes typically formed during its interval $I_{i}$, which is crucial due to the irreversibility of details, and (ii) avoid contradictions likely to emerge at that stage, leveraging the flexibility in high-frequency details.
+This decomposition conditions the diffusion model on contextually coherent content that evolves in tandem with the coarse-to-fine denoising process.
+To generate proxy prompts and their intervals, we use a large language model (LLM) that detects contextual contradictions and proposes suitable substitutes for conflicting concepts.
+It also infers the appropriate staging of these concepts within the proxy prompts.
+We implement this using a structured prompt template containing instructional text, in-context examples, and explanations of contextual contradictions.
+The examples demonstrate both successful decompositions and cases requiring no decomposition, enabling the LLM to generalize.
+The full instruction prompt is provided in the Appendix.
+In-context Examples.
+Our in-context examples take a target prompt as input and output proxy prompts with timestep intervals, along with a brief explanation of the contradiction.
+Requiring the LLM to provide this explanation encourages reasoning about conflicts and ensures coherent substitutions.
+These examples were created by identifying contextually contradictory prompts that fail under the base model (FLUX) and manually decomposing them into proxy prompts with corresponding intervals.
+The explanations were auto-generated by an LLM.
+We include 20 examples that demonstrate diverse strategies for handling contradictions.
+We next elaborate on one of the most frequent strategies in our method.
+Concept Substitution.
+In this strategy, a conflicting concept is temporarily replaced with a structurally appropriate proxy (Figure 4).
+A simpler alternative is to omit the conflicting concept, but introducing an object only in later stages without a placeholder can distort its perceived size or cause it to be omitted entirely.
+In Figure 6, we demonstrate this by comparing decompositions that differ only in when the second interval begins.
+The first proxy specifies the background, while the second adds the foreground.
+Introducing the foreground early allows the model to allocate more space, whereas delaying it constrains the layout and produces a smaller object.
+Substitution resolves these issues and yields stable layouts.
+In Figure 7, we show the effect of misplacing intervals across denoising stages.
+Using two proxies with different intervals, where the second is the full prompt, we observe two failure modes: introducing the full prompt too early prevents disentangling contradictions, while introducing it too late alters only fine details.
+Earlier in Figure 4 we illustrated the importance of careful proxy selection.
+4.3 Stage-aware Prompt Injection
+Given a sequence of proxy prompts $\{p_{1},p_{2},...,p_{n}\}$, their corresponding timestep intervals $\{I_{1},I_{2},...,I_{n}\}$, and a text-to-image (T2I) diffusion model, we condition the model using different prompts throughout the denoising process.
+At each timestep $t$, we apply the prompt $p_{i}$ such that $t\in I_{i}$.
+By aligning each proxy prompt with its interval, the denoising process is guided by concepts appropriate to the level of detail emerging at that stage.
+This enables gradual image construction while avoiding conflicts with the model’s learned priors.
+The injection mechanism integrates seamlessly into existing inference pipelines without architectural modifications and is compatible with a range of pretrained diffusion models.
+5 Experiments and Results
+In this section, we evaluate SAP through qualitative (Section 5.1) and quantitative (Section 5.2) comparisons.
+We further conduct ablation studies (Section 5.3) on component contributions and robustness, followed by a discussion on the limitations of our method.
 “A monkey is sitting on a white sandy beach.”
@@
 Step 1 Step 3 Step 5
-Figure 4.
-The effect of the object removal strategy.
-Each image is generated from a prompt that initially describes only the background (“a white sandy beach,” “a desert”).
-At the timestep indicated below each column, the foreground object (“a monkey,” “a snowman”) is added to the prompt.
-Removing the object from the initial prompt, without using a placeholder, significantly affects its apparent scale.
-Early insertion allows the model to allocate more spatial area, while late insertion confines the object to limited regions within the established layout.
-3.3. Stage-aware Prompt Injection
-Given a sequence of proxy prompts ${p_{1},p_{2},...,p_{n}}$, their corresponding timestep intervals ${I_{1},I_{2},...,I_{n}}$, and a text-to-image (T2I) diffusion model, we condition the model using different prompts throughout the denoising process.
-At each timestep $t$, we apply the prompt $p_{i}$ such that $t\in I_{i}$.
-By aligning each proxy prompt with a specific timestep interval, the denoising process is guided at each stage by concepts relevant to the level of detail emerging at that point.
-This enables the gradual construction of the image while avoiding conflicts with the model’s learned priors.
-This injection mechanism integrates seamlessly into existing inference pipelines without any architectural modifications, and is compatible with a range of pretrained diffusion models.
-Differences in denoising schedules across models are handled entirely during prompt decomposition (Section 3.2).
-We show examples with SD3 in the appendix (see Figure 11).
-“A camping tent is inside a bed room”
+Figure 6: Early insertion of the foreground allows the model to allocate more space to it, while late insertion confines the object within the existing layout, making it appear smaller (e.g., the snowman introduced at Step 5).
+Step labels indicate when the foreground object is introduced.
+Implementation Details.
+We use FLUX.1 [dev] [17] as the base T2I model and GPT-4o [1] for prompt decomposition.
+In all experiments, inference is performed using $T=50$ steps and the LLM is restricted to at most three proxies per prompt.
+Baseline hyperparameters follow their original papers or implementations, with $T=50$ steps for fair comparison.
+To further demonstrate robustness, we also report results with SD3.0 using the same LLM-generated proxy prompts and intervals.
+Baselines.
+We compare SAP against the following approaches: (1) base models FLUX-dev (denoted by FLUX) [17] and SD3.0 [11]; (2) R2F [22], a training-free method reported under three settings: SD3.0 (original), FLUX-schnell (official), and our reimplementation on FLUX; (3) Annealing Guidance [34], which trains a small MLP to predict the classifier-free guidance scale at each step; and (4) Ella [16], a fine-tuned model on SD1.5.
+(5) VL-DNP [8], a VLM-guided negative prompting method, evaluated using SD3 with Qwen2.5-VL-7B.
+Datasets
+We evaluate SAP using three datasets: Whoops!
+[7], Whoops-Hard, and ContraBench.
+Whoops!
+consists of 500 prompts paired with commonsense-defying images, designed to test visual reasoning and compositional understanding.
+While relevant to our task, many of its prompts are relatively easy for modern T2I models and do not consistently expose model limitations.
+To address this, we curate Whoops-Hard, a subset of 100 particularly difficult prompts from Whoops!, providing a stronger benchmark for evaluating current state-of-the-art models.
+To further probe semantic alignment under contradictory conditions, we introduce ContraBench, a curated set of 40 prompts capturing contextual contradictions.
+The dataset was constructed in two steps: (1) ChatGPT generated candidate prompts based on the definition of contextual contradictions, and (2) human annotators filtered them to retain only those that clearly expressed contradictions.
+The full prompt lists for both datasets are provided in the Appendix.
+5.1 Qualitative Results
+“A camping tent is inside a bedroom.”
 Refer to caption Refer to caption Refer to caption
@@
 Refer to caption Refer to caption Refer to caption
-too early too late Ours
-Figure 5.
-Impact of proxy object timing on generated image attributes.
-Changing the target object too early fails to disentangle the desired attribute, while modifying it too late prevents effective binding.
-Attribute control is most successful when applied within the correct time interval.
-In the top row, the initial proxy prompt is “A pillow fort in a bedroom”, while in the bottom row, the proxy prompt is “A mother duck guards three ducklings”.
-4. Experiments and Results
-In this section, we conduct qualitative (Section 4.1) and quantitative (Section 4.2) comparisons to assess the effectiveness of our method, along with ablation study (Section 4.3) to evaluate the contribution of individual components.
-We also present some limitations of our method (Section 4.4).
-Implementation Details.
-We use the state-of-the-art text-to-image model FLUX.1 [dev] (Labs, 2024) and GPT-4o as LLM to demonstrate the effectiveness of our approach.
-Inference is performed using $T=50$ sampling steps.
-In all experiments, we guide the LLM to limit the number of proxy prompts generated during decomposition to a maximum of three per input prompt.
-We use the baselines hyperparameters as appeared in their original papers or implementation.
-To ensure a fair comparison, we also use $T=50$ sampling steps for all baseline methods.
-Baselines
-Since our method builds upon FLUX.1 [dev], we directly compare our results against this base model to clearly demonstrate the improvements achieved by our approach.
-Additionally, we include a comparison to the training-free method R2F (Park et al., 2024), which similarly employs a large language model (LLM) to alternate between different prompts during generation.
-As R2F is implemented based on SD3 (Esser et al., 2024), we also compare our results to this model.
-Lastly, we compare against the fine-tuned model Ella (Hu et al., 2024), which is currently publicly available with SD 1.5 as its base model.
-Datasets
-We evaluate our method using three datasets: Whoops!
-(Bitton-Guetta et al., 2023), Whoops-Hard, and ContraBench.
-Whoops!
-consists of 500 prompts paired with commonsense-defying images, designed to test visual reasoning and compositional understanding.
-While relevant to our task, many of its prompts are relatively easy for modern T2I models and fail to consistently reveal their limitations.
-To address this, we curate a more challenging dataset Whoops-Hard, a subset comprising 100 particularly difficult prompts from Whoops!
-(Bitton-Guetta et al., 2023).
-This challenging benchmark provides a stronger basis for evaluating text-to-image generation given the current performance of state-of-the-art methods.
-Furthermore, to test more challenging cases of semantic alignment, we introduce a new benchmark ContraBench, a curated set of 40 prompts crafted to induce strong contextual contradictions, including background-object mismatches, semantic tension, and scale conflicts.
-ContraBench serves as a focused, high-difficulty benchmark for evaluating text-image alignment on prompts that contain contextual contradictions.
-We will make both ContraBench and the Whoops-Hard subset publicly available.
+Step 1 Step 14 Ours
+Figure 7: Effect of interval assignment.
+Introducing the full prompt too early fails to disentangle contextual contradictions, while introducing it too late alters only fine details.
+Top proxy: “A pillow fort in a bedroom”; bottom proxy “A mother duck guards three ducklings”.
 “A monkey juggles tiny elephants.”
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 “Shrek is blue.”
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 “Lightning striking a dilapidated shack on a clear sunny day.”
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Ella SD3 R2F FLUX Ours
-Figure 6.
-Our method effectively resolves contextual contradictions across diverse forms, including inconsistent backgrounds, unnatural color assignments, and implausible actions.
-Compared to prior approaches, our results exhibit higher semantic coherence and more accurate prompt grounding under contradictory prompts.
-4.1. Qualitative Results
-Comparisons with baselines.
-Figures 6 and 9 present qualitative comparisons on the Whoops!
-and ContraBench datasets.
-Across both benchmarks, the baseline methods FLUX, SD3, R2F, and Ella consistently exhibit characteristic failure modes when handling prompts with contextually contradictory concepts.
-For instance, as shown in Figure 6, our method is the only one to successfully generate scenarios such as a blue Shrek or a monkey juggling tiny elephants.
-For SD3 and FLUX, these prompts conflict with their learned priors, leading to failures in generating text-aligned images.
-Ella was originally designed for long and complex prompts, not specifically for those containing contextual contradictions, and therefore struggles with challenging contradiction cases.
-The training-free baseline R2F alternates between prompts at predefined timesteps, a strategy originally designed for attribute binding rather than addressing contextual contradictions.
-While it can reinforce rare concepts, this approach does not align prompt content with the stages at which semantic features emerge during denoising.
-As a result, it often produces hybrid concepts that merge incompatible elements from conflicting objects (see the owl and SpongeBob in Figure 9).
-In contrast, our method produces semantically coherent outputs by introducing proxy prompts at denoising stages where the corresponding visual features are expected to emerge.
-This alignment enables the model to effectively handle conflicting concepts.
-Across both Whoops!
-and ContraBench, our approach consistently generates contextually contradictory elements with high accuracy.
-To demonstrate the robustness of our approach, we provide a comparison across several random seeds in Figure 10.
-The same seeds are used for all methods, showing that our method consistently outperforms the baselines.
-Improve realism
-Figure 7 demonstrates how our method generates photorealistic and semantically coherent images for prompts with atypical attribute combinations.
-In contrast, FLUX often defaults to cartoon-like renderings, likely due to priors associating fantastical elements with imaginative visual styles.
-By using contradiction-free proxy prompts, our method enables the model to produce realistic outputs without explicitly requiring photorealism in the prompt.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Ella Annealing SD3 R2F FLUX R2F (schnell) R2F (FLUX) SAP
+Figure 8: Qualitative comparison of SAP with baseline methods.
+Our method resolves contextual contradictions, whereas baselines struggle to produce text-aligned images.
+Additional examples are provided in the Appendix.
+Figure 8 presents qualitative comparisons on the Whoops!
+and ContraBench datasets (see Appendix for more results).
+Across both benchmarks, baseline methods consistently exhibit characteristic failure modes when handling contradictory prompts.
+In contrast, SAP successfully generates challenging cases such as a blue Shrek or a monkey juggling tiny elephants (Figure 8).
+In addition to FLUX, SAP also improves SD3 generations under contradictory prompts (Figure 9).
+For SD3 and FLUX, contradictory prompts expose conflicts with learned priors, resulting in prompt misalignment.
+Ella and Annealing Guidance, not designed for contradictions, perform less effectively on such cases.
+R2F alternates between prompts at predefined timesteps, a strategy designed for attribute binding rather than addressing contextual contradictions.
+While it can reinforce rare concepts, it does not align prompts with the stages at where semantic features emerge during denoising.
+As a result, it often produces hybrid concepts that merge incompatible elements from conflicting concept (see the bodybuilder in Figure 9, and the owl and SpongeBob results in the Appendix).
+In contrast, SAP produces semantically coherent outputs by introducing proxy prompts at denoising stages where corresponding features emerge.
+This enables effective handling of conflicting concepts.
+Across both benchmarks, SAP consistently generates coherent images for contradictory prompts.
+Robustness is further demonstrated in the Appendix, comparing generations across multiple seeds.
+5.2 Quantitative Results
+We evaluate prompt alignment using GPT-4o vision–language model (VLM).
+For each generated image, GPT-4o assigns a score from 1–5 based on adherence to the prompt.
+Scores are averaged across three fixed random seeds per prompt and scaled to 20–100.
+The evaluation prompt is provided in the Appendix.
+As shown in Table 1, SAP outperforms all baselines across the three benchmarks.
+Between the base models, SD3.0 tends to yield stronger alignment under contradictions, while FLUX offers higher visual quality (Figure 8).
+SAP improves both backbones, enhancing prompt adherence while maintaining the visual fidelity of the underlying models (Figures 8 and 9 ).
+SD3 $\text{R2F}_{\text{SD3}}$ $\textbf{SAP}_{\text{SD3}}$ Refer to caption Refer to caption Refer to caption “A grown man has a baby’s pacifier in his mouth.” Refer to caption Refer to caption Refer to caption “A bodybuilder balancing on pointe shoes.”
+Figure 9: SAP is robust to the base model, as shown by the results obtained with $\text{SAP}_{\text{SD3}}$.
+User study.
+VLM-based metrics often miss subtle semantic inconsistencies and do not adequately assess image quality.
+To complement them, we conducted a user study evaluating both prompt adherence and overall visual appearance.
+We randomly sampled 24 prompts from the Whoops!
+and ContraBench benchmarks.
+For each prompt, participants compared two images, one generated by SAP and the other by a baseline, and answered: (1) which most accurately reflected the prompt, and (2) which had higher visual quality.
+In total, we collected responses from 61 users, yielding 1,464 individual evaluations.
+Table 2 summarizes the win rates of SAP against each baseline.
+These results highlight the superiority of our approach in handling contextually contradictory prompts, achieving both stronger prompt alignment and higher visual quality.
+5.3 Ablation Studies
+We evaluate SAP through ablations that assess design choices in prompt decomposition and robustness under different conditions, such as timestep perturbations.
+Additional results on alternative LLMs and non-contradictory prompts are in the Appendix.
+Table 1: Quantitative evaluation on various benchmarks using the GPT-4o vision-language model.
+We report average alignment, where alignment reflects how well the image matches the prompt semantics, independent of visual quality.
+SAP achieves the best results, regardless of the base model.
+Benchmarks
+Models Whoops Whoops- Contra-
+Hard Bench
+SD3.0 82.63 55.73 57.5
+FLUX 78.85 44.3 57.16
+Ella 69.09 45.19 55.16
+Annealing 79.59 59.06 58.33
+VL-DNP 82.79 56.26 58.83
+R2F 83.50 57.06 59.16
+$\text{R2F}_{\text{schnell}}$ 79.58 54.80 59.33
+$\text{R2F}_{\text{FLUX}}$ 48.68 32.80 25.33
+$\text{SAP}_{\text{SD3.0}}$ 85.87 64.40 65.33
+SAP 85.10 62.13 66.16
+Table 2: User study results. Win rates of SAP in text-image alignment and image quality, compared against each baseline method.
+SD3 FLUX Ella Annealing R2F
+Alignment $70\%$ $\% 8181\%$ $73\%$ $75\%$
+Quality $72\%$ $63\%$ $74\%$ $79\%$ $68\%$
+Table 3: Ablation on Whoops-Hard.
+We evaluate our prompt decomposition method by (1) removing in-context examples, (2) removing the explanation requirement, and (3) limiting decomposition to two proxy prompts.
+static w/o in-context w/o reasoning 2 proxy Full
+Alignment 44.3 48.0 46.46 60.26 62.13
+Prompt decomposition components.
+We conduct ablations on the Whoops-Hard benchmark, where each variant isolates a design choice to quantify its effect on alignment within our method (Table 3).
+In-context examples significantly improve the LLM’s ability to decompose contradictory prompts, leading to better text–image alignment.
+Removing the explanation requirement impairs reasoning and causes a notable drop, showing that generating explicit explanations encourages more coherent semantic decisions.
+Restricting decomposition to two proxies performs close to the full method, while allowing up to three proxies provides extra flexibility for harder cases and yields further gains.
+Robustness to LLM-predicted timestep intervals.
+Our method relies on LLM-predicted intervals to schedule proxy prompts, but these boundaries do not require exact placement.
+The earlier results (Figure 7) highlight that placing proxy prompts at the wrong stage of denoising (e.g., too early or too late) can harm alignment.
+Here we show that within the correct coarse stage, the method is robust to moderate boundary shifts.
+Specifically, we perturb interval boundaries while keeping the proxy prompts fixed, uniformly shifting them within windows of varying size.
+As shown in Table 4, small shifts of up to $\pm 1$ step (window=3) have almost no effect on alignment, and even larger shifts of up to $\pm 2$ steps (window=5) cause only minor degradation, despite representing a substantial perturbation relative to the full method effective range ($\sim$12 steps; see Appendix ).
+These results confirm that SAP is sensitive to the stage at which information is introduced, but largely insensitive to exact step boundaries within that stage.
+Refer to caption Refer to caption Refer to caption
+“A bouquet of flowers is upside down in a vase” “A white glove has 6 fingers” “The shadow of a cat is facing the opposite direction”
+Figure 10: Failure cases of SAP.
+Limitations.
+In Figure 10, we present representative failure cases of our method on examples from the Whoops!
+benchmark.
+Since our approach relies on guiding the model through text alone, it cannot control properties that the underlying model inherently struggles with, such as generating specific quantities or enforcing precise orientations; a more detailed categorization and analysis of failure modes is provided in the supplementary.
+Table 4: Effect of perturbing LLM-predicted timestep intervals. Boundaries are uniformly shifted within window $w$. $SAP_{w=i}$ denotes evaluation with window ii.
+Benchmarks
+Models Whoops Whoops- Contra-
+Hard Bench
+FLUX 78.85 44.3 57.16
+SAP 85.10 62.13 66.16
+$\text{SAP}_{w=3}$ 84.18 62.06 65.5
+$\text{SAP}_{w=5}$ 81.46 58.46 62.5
+6 Conclusions
+We introduced a training-free framework for resolving contextual contradictions in text-to-image generation, cases where seemingly plausible prompts fail due to strong, hidden model biases.
+At its core, our method leverages the coarse-to-fine generation process to separate contradictions across denoising stages, enabling faithful rendering of prompts that would otherwise yield semantically inconsistent outputs.
+The introduction of proxy prompts steers the generative process in line with the model’s priors, enabling it to resolve conflicts and preserve semantic fidelity without the need for retraining or fine-tuning.
+We argue that since our approach already leverages the broad world knowledge of vision–language models, integrating them more tightly with generative models holds promise for addressing contextual contradictions directly.
+As a next step, we plan to explore emerging compound architectures that combine VLMs and generative models, with the aim of understanding how to effectively harness them to resolve conflicts in open-ended generation.
+Acknowledgments
+This research was supported in part by the Israel Science Foundation (grants no. 2492/20 and 1473/24), Len Blavatnik and the Blavatnik family foundation.
+Supplementary Material
+7 Ethics Statement
+Our work contributes to improving the semantic alignment of text-to-image models under contradictory or biased prompts.
+As a consequence, our method enhances users’ ability to control generative models and faithfully render contradictory concepts.
+While this provides positive benefits, such as reducing unintended biases and enabling more inclusive image generation, it also increases the potential for misuse, including the creation of harmful, misleading, or inappropriate content.
+As with any advance in generative modeling, these dual-use concerns highlight the importance of responsible deployment, safeguards, and continued ethical oversight to ensure that such improvements contribute positively to society.
+8 Additional Results
+Robustness across LLMs.
+Since our framework hinges on LLM-driven prompt decomposition, we further examined its robustness under different language models.
+We evaluated both a proprietary model (GPT-4o) and a comparatively lightweight open-source alternative (LLaMA-3.1-8B-Instruct).
+While GPT-4o delivers the strongest performance, the smaller LLaMA-3.1-8B-Instruct still yields consistent improvements over the baseline (see Table 5).
+Improved realism.
+SAP generates photorealistic and semantically coherent images for prompts with atypical attribute combinations (Figure 11).
+In contrast, FLUX often defaults to cartoon-like renderings, even when photorealism is explicitly requested, revealing a contextual contradiction between fantastical content and realistic style.
+By using contradiction-free proxy prompts, SAP avoids these biases and produces realistic outputs regardless of whether photorealism is explicitly required in the prompt.
+Non-contradictory prompts.
+To ensure applicability in general text-to-image scenarios, we verify that our method does not negatively affect prompts without contextual contradictions.
+We find that including even a single non-contradictory in-context example is sufficient for the LLM to default to using the full prompt in such cases.
+We evaluate this behavior using GPT-4o alignment scores on the PartiPrompts-Simple benchmark, which contains simple, non-contradictory prompts (Table 6).
+Additional qualitative comparisons.
+Figures 13 and 14 present additional qualitative comparisons of our method, while Figure 15 shows results across multiple seeds.
+Table 5: Performance of SAP when combined with different LLMs, comparing GPT-4o and Llama-3.1-8B-Instruct.
+Benchmarks
+Models Whoops Whoops- Contra-
+Hard Bench
+FLUX 78.85 44.3 57.16
+$\text{SAP}_{\text{GPT4o}}$ 85.10 62.13 66.16
+$\text{SAP}_{\text{Llama3.1}}$ 80.52 59.53 61.16
+Table 6: Alignment performance on the PartiPrompts-simple benchmark, which contains simple, non-contradictory prompts.
+Scores are computed using GPT-4o vision-language model.
+Our method achieves comparable performance to the base model, indicating no degradation on regular prompts.
+Models PartiPrompts-simple
+FLUX 93.46
+SAP 93.06
 “A kangaroo delivering mail on a bicycle.” “A boy is kissing a hedgehog.”
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
-Ours
-Figure 7.
-FLUX tends to generate realistic images by default.
+SAP
+Figure 11: FLUX tends to generate realistic images by default.
 However, when given unrealistic prompts, it often produces cartoon-like samples.
 In contrast, our method, which gradually resolves such prompts through coherent proxy stages, consistently generates realistic and semantically aligned images.
-4.2. Quantitative Results
-Our method focuses on solving alignment with challenging and contradictory textual prompts.
-Therefore, we evaluate its performance by measuring prompt alignment.
-To do so, we use the GPT-4o vision-language model (VLM).
-The model is prompted to assign a score from 1 to 5 for each generated image, based on how well the image aligns with the given textual prompt.
-The exact evaluation prompt is provided in the appendix (See Table 6).
-For each prompt, we generate outputs using three fixed random seeds, which are applied consistently across all evaluated methods.
-The average score across the three seeds is then scaled to a range between 20 and 100 by multiplying by 20.
-As can be seen in Table 1, our method outperforms the baselines across all three benchmarks.
-Notably, using SD3 as the base model yields better prompt alignment than the state-of-the-art FLUX model under contradictory prompt settings, while FLUX demonstrates superior visual quality (See Figure 9).
-In contrast, our method, built on top of FLUX, achieves the highest prompt adherence while preserving its strong visual quality.
-This results in the best overall performance, combining high-fidelity generation with robust alignment to challenging prompts.
-Table 1.
-Quantitative evaluation on the Whoops dataset using the GPT-4o vision-language model.
-We report average alignment, where alignment reflects how well the image matches the prompt semantics, without looking at the quality of the image
-Benchmarks
-Models Whoops Whoops-Hard ContraBench
-Ella 69.09 45.19 55.16
-SD3.0 82.63 55.73 57.5
-R2F (SD3.0) 83.50 57.06 59.16
-FLUX 78.85 44.3 57.16
-Ours (FLUX) 85.10 62.13 66.16
-Table 2. User study results. Win rates of our method in text-image alignment and image quality, compared against each baseline method.
-Ella SD3.0 R2F (SD3.0) FLUX
-Alignment $81\%$ $70\%$ $75\%$ $81\%$
-Quality $74\%$ $72\%$ $68\%$ $63\%$
-User study
-VLM-based automatic metrics, such as those shown in Table 1, often fail to detect subtle semantic inconsistencies, particularly when evaluating generations conditioned on complex or contradictory prompts.
-Moreover, these models do not adequately assess the visual quality of the generated images.
-To address these limitations, we conducted a user study aimed at evaluating both prompt adherence and visual quality from a human perspective.
-We randomly sampled 24 prompts from the Whoops!
-and ContraBench benchmarks.
-For each prompt, participants were shown two images: one generated by our method and the other by one of the baselines.
-They were asked to answer two questions: (1) which image most accurately reflected the prompt, and (2) which image had better visual quality.
-In total, we collected responses from 61 users, resulting in 1,464 individual evaluations.
-The results of the user study are summarized in Table 2, which reports the win rate of our method when compared pairwise against each baseline.
-These results highlight the superiority of our approach in handling contextually contradictory prompts, achieving both better prompt alignment and visual quality compared to baseline approaches.
-Refer to caption Refer to caption Refer to caption
-“A bouquet of flowers is upside down in a vase” “A white glove has 6 fingers” “The shadow of a cat is facing the opposite direction”
-Figure 8. Failure cases from Whoops!.
-4.3. Ablation Studies
-To evaluate the contribution of each component in our method, we perform an ablation study on the Whoops-Hard benchmark by selectively removing or modifying individual elements of the method, and present the results in Table 3.
-Each configuration isolates a specific design choice to quantify its effect on alignment performance.
-We observe that using in-context examples significantly improves the LLM’s ability to decompose and interpret contextually contradictory prompts, resulting in better text-image alignment.
-Removing the explanation requirement from the LLM impairs its reasoning capabilities and leads to a notable performance drop, suggesting that requiring the LLM to generate explicit explanations helps it make more coherent semantic decisions, ultimately improving prompt decomposition.
-Additionally, using only two proxy prompts, underperforms compared to the full method, highlighting the importance of introducing prompt information at targeted stages for robust alignment.
-4.4. Limitations
-In Figure 8, we present three failure cases of our method on examples from the Whoops!
-benchmark.
-Since our approach relies on guiding the model through text alone, it cannot control properties that the underlying model inherently struggles with—such as generating specific quantities or enforcing precise orientations.
-Table 3.
-Ablation study on the Whoops-Hard Benchmark.
-We evaluate the impact of key components of our prompt decomposition method by: (1) removing in-context examples, (2) removing the explanation requirement, and (3) restricting the decomposition to at most two proxy prompts.
-Results show a significant drop in alignment when any of these components are removed.
-static w/o in-context w/o reasoning 2 proxy Full
-Alignment 44.3 48.0 46.46 50.73 62.13
-5. Conclusions
-We have introduced a training-free framework for resolving contextual contradictions in text-to-image generation – that is, cases where seemingly plausible prompts are internally inconsistent because of strong, hidden biases in the model.
-Our method identifies and strategically separates such contradictions over time, enabling faithful rendering of prompts that would otherwise yield semantically inconsistent outputs.
-At its core, our approach leverages the broad real-world understanding embedded in large language models to reason about the semantic plausibility of prompts.
-Rather than modifying the training data or fine-tuning the model, we harness this reasoning capacity at inference time to decompose prompts into a sequence of contextually coherent stages.
-These proxy prompts guide the generative process in alignment with the model’s prior expectations, helping it circumvent conflicts and maintain semantic fidelity.
-Extensive experiments demonstrate that our stage-aware decomposition substantially improves alignment and image quality – especially for prompts that challenge the model’s default associations.
-Future directions include refining our method using VLM-generated responses to improve proxy prompts generation dynamically during the generation process, and applying our method to text-to-video generation models.
-Appendix A Appendix
-Table 4. Full LLM prompt instruction for time step dependent prompt conditioning, used to decompose prompts by denoising stage.
+Failure Mode Analysis.
+We analyze failure cases of SAP and group them into three categories, summarized in Table 7.
+The majority of failures arise from incorrect proxy prompt assignment, where the selected proxy does not adequately preserve the intended structure or semantics of the target concept.
+In many such cases, manually correcting the proxy resolves the contradiction, as illustrated in Figure 12 (left), indicating that these failures stem from limitations of the LLM rather than the proposed framework, and can potentially be mitigated.
+A second group of failures stems from inherent limitations of the underlying diffusion model, such as difficulty generating an empty pool or rendering unusual symbolic structures (e.g., compass directions), which persist even with correct prompt decomposition (Figure 12, right).
+Finally, only a small fraction of failures is attributed to suboptimal interval assignment, consistent with our findings that SAP is robust to moderate variations in timestep scheduling.
+Table 7: Breakdown of SAP failure cases by root cause.
+Interval Proxy Model
+assignment assignment limitation
+Failure cases 3.3% 63.3% 33.3%
+Bad proxy / Corrected Model limitation
+Refer to caption Refer to caption Refer to caption Refer to caption
+“A man is welding
+without a mask”
+“A man jumping
+into an empty
+swimming pool.”
+“A compass with
+North South South
+West points”
+Figure 12: Failure cases of SAP.
+Left: a proxy assignment failure, where an unsuitable proxy leads to incorrect generation; manually correcting the proxy (right image in pair) resolves the contradiction.
+Right: failures due to inherent limitations of the underlying diffusion model, such as difficulty generating an empty pool or rendering unusual symbolic structures (e.g., a compass with incorrect directions).
+Computational Overhead.
+SAP introduces modest overhead due to a single LLM inference for prompt decomposition, performed once per prompt prior to diffusion, and a lightweight additional embedding pass.
+As this step occurs only once and is independent of the denoising process, it does not significantly impact overall generation time.
+9 LLM Instruction for Prompt Decomposition
+Tables 8 and 9 detail the full LLM instruction used for our method’s decomposition, along with the corresponding in-context examples.
+In a single inference pass, our method detects contextual contradictions, generates proxy prompts, and assigns timestep intervals.
+Table 10 presents examples of our LLM input prompts, along with the corresponding output explanations and the decomposition into proxy prompts and timestep intervals.
+10 Provided Benchmarks
+We describe the construction of ContraBench and Whoops-Hard in the main text. Here, we provide the full lists of prompts for these benchmarks in Table 11 and Table 12, respectively.
+11 VLM Evaluation
+We utilize GPT-4o to assess alignment between prompts and their generated images. The instruction prompt provided to the VLM is shown in Table 13.
+Ella Annealing SD3 R2F Flux SAP
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A balloon is lifting up a package.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A coocoo clock with a horse popping out.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“An owl is perched upside down on a branch.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A photorealistic photo of SpongeBob SquarePants dancing ballet.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“Children in Africa are building a snowman.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A cruise ship parked in a bathtub.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A fortune teller predicts the future with a basketball.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A Viking warrior flying an airplane.”
+Figure 13: Qualitative comparison. Our method consistently generates text-aligned images for contextually contradicting prompts.
+Ella Annealing SD3 R2F Flux SAP
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“An Egyptian pyramid is constructed using modern construction gear.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A cat balancing a skyscraper on its nose.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A cowboy swimming competitively in an Olympic pool.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A knight in chess is a unicorn.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A man giving a piggyback ride to an elephant.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A butterfly is in a bee’s hive.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“An eagle is swimming under-water.”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A photorealistic image of a bear ironing clothes in a laundry room”
+Figure 14: Qualitative comparison. Our method consistently generates text-aligned images for contextually contradicting prompts.
+Ella Annealing SD3 R2F Flux SAP
+Seed 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Seed 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Seed 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“a professional boxer does a split.”
+Seed 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Seed 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Seed 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A woman writing with a dart.”
+Seed 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Seed 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Seed 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A man riding a jet ski through the desert.”
+Figure 15: Qualitative comparison across multiple seeds. Our method consistently generates text-aligned images for contextually contradicting prompts.
+Table 8: Full LLM prompt instruction SAP, used to decompose prompts by denoising stages.
 <System Prompt>
@@
   "switch_prompts_steps": [<step1>, <step2>, ...]
-- The length of switch_ prompts_steps should be one less than prompts_list. - Do not include any text outside this structure.
-A.1. LLM Instruction for Prompt Decomposition
-Tables 4 and 5 detail the full LLM instruction used for our method’s decomposition, along with the corresponding in-context examples.
-In a single inference pass, our method detects contextual contradictions, generates proxy prompts, and assigns timestep intervals.
-Table 8 presents examples of our LLM input prompts, along with the corresponding output explanations and the decomposition into proxy prompts and timestep intervals.
-Table 5. In-context examples for time step dependent prompt conditioning, illustrating reasoning steps and prompt decomposition across denoising stages.
+- The length of switch_ prompts_steps should be one less than prompts_list.
+- Do not include any text outside this structure.
+Table 9: In-context examples SAP, illustrating reasoning steps and prompt decomposition across denoising stages.
 Input: A polar bear in a desert
@@
   "switch_prompts_steps": []
-A.2. VLM evaluation
-We evaluate text-to-image alignment for the different models on the benchmarks using the VLM GPT-4o. The instruction prompt used for the VLM is presented in Table 6.
-A.3. ContraBench
-To evaluate image generation under prompts containing contextual contradictions, we introduce ContraBench, presented in Table 7.
-This dataset consists of 40 prompts generated using ChatGPT, each intentionally constructed to include visual or semantic inconsistencies challenge standard diffusion models.
-The prompts span a range of contradiction types, including physical implausibility, role violation, object misplacement, and surreal transformations.
-ContraBench serves as a targeted benchmark for assessing a model’s ability to produce visually coherent and semantically plausible images in the presence of contradictory concepts.
-A.4. Method Generalization to Different Models
-Our method’s proxy prompt injection mechanism integrates seamlessly into existing inference pipelines without requiring any architectural modifications, and it is compatible with a range of pretrained diffusion models.
-Differences in denoising schedules across models are handled entirely during the prompt decomposition stage.
-See Figure 11 for an example of how our method integrates with SD3 without any modifications.
-Table 6. VLM instruction for prompt alignment. Instructions provided to the VLM for scoring semantic alignment of generated images.
+Table 10: Prompt decompositions of contradictory prompts using SAP.
+Original Prompt Explanation Prompt Sequence Switch Step(s)
+A grown man has a baby’s pacifier in his mouth.
+The grown man is established first for structure, while the pacifier, being a small but semantically significant detail, is added once the main identity is resolved.
+1.
+A grown man with a small object in his mouth 2.
+A grown man has a baby’s pacifier in his mouth 4
+A dragon is blowing water.
+Dragons are more commonly depicted blowing fire.
+A proxy of white smoke is visually similar to water mist in texture, stabilizing the emission process before resolving the surreal water emission.
+1.
+A dragon blowing white smoke 2.
+A dragon blowing water 3
+A pizza with grape toppings.
+Pizza with traditional toppings stabilizes the geometry and color before introducing the visually similar yet unusual grape topping.
+1.
+A pizza with pepperoni toppings 2.
+A pizza with grape toppings 3
+A coin floats on the surface of the water.
+Coins typically sink in water, not float.
+Starting with a leaf—an object that naturally floats—ensures that this behavior within the scene is handled correctly before introducing the coin.
+1.
+A leaf floats on the surface of the water 2.
+A coin floats on the surface of the water 4
+A cockatoo parrot swimming in the ocean.
+Cockatoos are birds and naturally do not swim; starting with a simple bird on water stabilizes position and motion.
+Progressing to a duck, before introducing the cockatoo parrot, eases the transition into the final surreal visual.
+1.
+A duck swimming in the ocean 2.
+A parrot swimming in the ocean 3.
+A cockatoo parrot swimming in the ocean 3, 6
+Shrek is blue.
+Shrek is a distinct character with a recognizable green color.
+Using a simple b̈lue ogreïnitially sets the stage for a color change before fully introducing Shrek to ensure visual coherence.
+1.
+A blue ogre 2.
+Shrek is blue 3
+A professional boxer does a split.
+Professional boxers are typically shown in athletic stances related to fighting, not performing a split.
+Starting with a gymnast performing a split supports the action, introducing a boxer in similar attire balances identity shift without disrupting the pose.
+1.
+A gymnast performing a split 2.
+A boxer performing a split 3.
+A professional boxer doing a split 3, 6
+Table 11: ContraBench. A curated benchmark of 40 contradictory prompts for evaluating text-to-image models.
+ID Prompt ID Prompt
+1 A professional boxer does a split 21 A mosquito pulling a royal carriage through Times Square
+2 A bear performing a handstand in the park 22 A grandma is ice skating on the roof
+3 A bodybuilder balancing on point shoes 23 A baseball player backswing a yellow ball with a golf club
+4 A chicken is smiling 24 A house with a circular door
+5 A cruise ship parked in a bathtub 25 A photorealistic image of a bear ironing clothes in a laundry room
+6 A man giving a piggyback ride to an elephant 26 A pizza being used as an umbrella in the rain
+7 A zebra climbing a tree 27 A cubist lion hiding in a photorealistic jungle
+8 A coffee machine dispensing glitter 28 A cowboy swimming competitively in an Olympic pool
+9 A vending machine in a human running posture 29 A realistic photo of an elephant wearing slippers
+10 A ballerina aggressively flipping a table 30 A computer mouse eating a piece of cheese
+11 A bathtub floating above a desert in a tornado 31 A horse taking a selfie with a smartphone
+12 A monkey juggles tiny elephants 32 A sheep practicing yoga on a mat
+13 A woman has a marine haircut 33 A snake eating a small golden guitar
+14 A tower with two hands 34 A soccer field painted on a grain of rice
+15 An archer is shooting flowers with a bow 35 A snake with feet
+16 A muscular ferret in the woods 36 A woman brushing her teeth with a paintbrush
+17 A barn built atop a skyscraper rooftop 37 A horse with a hump
+18 A cat balancing a skyscraper on its nose 38 A hyperrealistic unicorn made of origami
+19 A cow grazing on a city rooftop 39 A library printed on a butterfly’s wings
+20 A fireplace burning inside an igloo 40 A photorealistic photo of SpongeBob SquarePants dancing ballet
+Table 12: Whoops-Hard. A curated subset of 100 challenging prompts from the Whoops! benchmark.
+ID Prompt ID Prompt
+1 A bouquet of flowers is upside down in a vase 51 A Japanese tea ceremony uses coffee instead of tea
+2 A man is welding without a mask 52 A wagon is being pushed from behind by two opposite facing horses
+3 A man eats hamburgers in a baby chair 53 The Girl with a Pearl Earring wears a golden hoop earring
+4 A turn right street sign with a left turn arrow 54 A chandelier is hanging low to the ground
+5 Goldilocks sleeps with four bears 55 The portrait of the Mona Lisa depicts a stern male face
+6 A cake wishes a happy 202nd birthday 56 A car with the steering wheel right in the middle of the dashboard
+7 Children are unhappy at Disneyland 57 A pagoda sits in front of the Eiffel Tower
+8 An orange carved as a Jack O’Lantern 58 A man without protection next to a swarm of bees
+9 A pen is being sharpened in a pencil sharpener 59 A kiwi bird in a green bamboo forest
+10 Steve Jobs demonstrating a Microsoft tablet 60 The Sphinx is decorated like a sarcophagus outside a Mayan temple
+11 Shrek is blue 61 A butterfly is in a bee’s hive
+12 A MacBook with a pear logo on it 62 A rainbow colored tank
+13 A woman hits an eight ball with a racket 63 Movie goers nibble on vegetables instead of popcorn
+14 Vikings ride on public transportation 64 A grown man has a baby’s pacifier in his mouth
+15 A gift wrapped junked car 65 A full pepper shaker turned upside down with nothing coming out
+16 A rainbow is filling the stormy sky at night 66 The Tiger King, Joe Exotic, poses with an adult saber-tooth tiger
+17 John Lennon using a MacBook 67 A scale is balanced with one side full and the other empty
+18 Michelangelo’s David is covered by a fig leaf 68 A pizza box is full of sushi
+19 Chuck Norris struggles to lift weights 69 A man wearing a dog recovery cone collar while staring at his dog
+20 Paratroopers deploy out of hot air balloons 70 A woman’s mirror reflection is wearing different clothes
+21 A train on asphalt 71 A woman using an umbrella made of fishnet in the rain
+22 Lionel Messi playing tennis 72 A field of sunflowers with pink petals
+23 A man jumping into an empty swimming pool 73 An eagle swimming under water
+24 An airplane inside a small car garage 74 A woman stands in front of a reversed reflection in a mirror
+25 An upside down knife about to slice a tomato 75 Stars visible in the sky with a bright afternoon sun
+26 Dirty dishes in a bathroom sink 76 A car with an upside down Mercedes-Benz logo
+27 A roulette wheel used as a dart board 77 An owl perched upside down on a branch
+28 A smartphone plugged into a typewriter 78 A man in a wheelchair ascends steps
+29 A passenger plane parked in a parking lot 79 Bach using sound mixing equipment
+30 Guests are laughing at a funeral 80 A steam train on a track twisted like a roller coaster
+31 A cat chasing a dog down the street 81 Roman centurions fire a cannon
+32 The Statue of Liberty is holding a sword 82 A crab with four claws
+33 A Rubik’s cube with ten purple squares 83 Elon Musk wearing a shirt with a Meta logo
+34 A girl roller skating on an ice rink 84 A compass with North South South West points
+35 A butterfly swimming under the ocean 85 A glass carafe upside down with contents not pouring
+36 Lightning striking a shack on a sunny day 86 Princess Diana standing in front of her grown son, Prince Harry
+37 The Cookie Monster is eating apples 87 A children’s playground set in the color black
+38 A man is given a purple blood transfusion 88 A mug of hot tea with a plastic straw
+39 An unpeeled banana in a blender 89 A whole pomegranate inside a corked glass bottle
+40 A square apple 90 Belle from Beauty and the Beast about to kiss the Frog Prince
+41 A place setting has two knives 91 A person’s feet facing opposite directions
+42 A koala in an Asian landscape 92 A bowl of cereal in water
+43 A mouse eats a snake 93 A boy playing frisbee with a porcelain disk
+44 A field of carrots growing above ground 94 A chef prepares a painting
+45 A pregnant woman eating raw salmon 95 A dragon blowing water
+46 A tiger staring at zebras in the savanna 96 The lip of a pitcher on the same side as the handle
+47 Albert Einstein driving a drag racing car 97 Greta Thunberg holding a disposable plastic cup
+48 A soccer player about to kick a bowling ball 98 A fortune teller predicting the future with a basketball
+49 An old man riding a unicycle 99 A balloon lifting up a package
+50 A hockey player drives a golf ball down the ice 100 Bruce Lee in a yellow leotard and tutu practicing ballet
+Table 13: VLM instruction for evaluation. Used by GPT-4o to score semantic alignment of generated images.
 You are an assistant evaluating an image on how well it aligns with the meaning of a given text prompt.
@@
 ### ALIGNMENT EXPLANATION: <explanation>
-SD3 Ours(SD3) SD3 Ours(SD3)
-Refer to caption Refer to caption Refer to caption Refer to caption
-“A grown man has a baby’s pacifier in his mouth.” “A camping tent is inside a bed room.”
-Refer to caption Refer to caption Refer to caption Refer to caption
-“A mouse eats a snake.” “A dragon is blowing water.”
-Figure 11. Generated images from our method applied on SD3.
-Table 7. ContraBench prompt set.
-ID Prompt
-1 A professional boxer does a split
-2 A bear performing a handstand in the park
-3 A photorealistic photo of SpongeBob SquarePants dancing ballet
-4 A cowboy swimming competitively in an Olympic pool
-5 A cruise ship parked in a bathtub
-6 A man giving a piggyback ride to an elephant
-7 A zebra climbing a tree
-8 A coffee machine dispensing glitter
-9 A vending machine in a human running posture
-10 A ballerina aggressively flipping a table
-11 A bathtub floating above a desert in a tornado
-12 A monkey juggles tiny elephants
-13 a woman has a marine haircut
-14 A tower with two hands
-15 An archer is shooting flowers with a bow
-16 A baseball player backswing a yellow ball with a golf club
-17 A barn built atop a skyscraper rooftop
-18 A cat balancing a skyscraper on its nose
-19 A cow grazing on a city rooftop
-20 A fireplace burning inside an igloo
-21 A mosquito pulling a royal carriage through Times Square
-22 A grandma is ice skating on the roof
-23 A muscular ferret in the woods
-24 A house with a circular door
-25 A photorealistic image of a bear ironing clothes in a laundry room
-26 A pizza being used as an umbrella in the rain
-27 A cubist lion hiding in a photorealistic jungle
-28 A chicken is smiling
-29 A realistic photo of an elephant wearing slippers
-30 A computer mouse eating a piece of cheese
-31 A horse taking a selfie with a smartphone
-32 A sheep practicing yoga on a mat
-33 a snake eating a small golden guitar
-34 A soccer field painted on a grain of rice
-35 A snake with feet
-36 A woman brushing her teeth with a paintbrush
-37 A horse with a hump
-38 A hyperrealistic unicorn made of origami
-39 A library printed on a butterfly’s wings
-40 A bodybuilder balancing on pointe shoes
-Table 8. Examples of prompt decomposition for contradictory prompts using our method.
-Original Prompt Explanation Prompt Sequence Switch Step(s)
-A grown man has a baby’s pacifier in his mouth.
-The grown man is established first for structure, while the pacifier, being a small but semantically significant detail, is added once the main identity is resolved.
-(1) A grown man with a small object in his mouth (2) A grown man has a baby’s pacifier in his mouth 4
-A dragon is blowing water.
-Dragons are more commonly depicted blowing fire.
-A proxy of white smoke is visually similar to water mist in texture, stabilizing the emission process before resolving the surreal water emission.
-(1) A dragon blowing white smoke (2) A dragon blowing water 3
-A pizza with grape toppings.
-Pizza with traditional toppings stabilizes the geometry and color before introducing the visually similar yet unusual grape topping.
-(1) A pizza with pepperoni toppings (2) A pizza with grape toppings 3
-A coin floats on the surface of the water.
-Coins typically sink in water, not float.
-Starting with a leaf—an object that naturally floats—ensures that this behavior within the scene is handled correctly before introducing the coin.
-(1) A leaf floats on the surface of the water (2) A coin floats on the surface of the water 4
-A cockatoo parrot swimming in the ocean.
-Cockatoos are birds and naturally do not swim; starting with a simple bird on water stabilizes position and motion.
-Progressing to a duck, before introducing the cockatoo parrot, eases the transition into the final surreal visual.
-(1) A duck swimming in the ocean (2) A parrot swimming in the ocean (3) A cockatoo parrot swimming in the ocean 3, 6
-A woman writing with a dart.
-Writing is typically associated with a pen or pencil.
-Starting with a pen stabilizes the action before introducing the unusual tool.
-(1) A woman writing with a pen (2) A woman writing with a dart 3
-Shrek is blue.
-Shrek is a distinct character with a recognizable green color.
-Using a simple b̈lue ogreïnitially sets the stage for a color change before fully introducing Shrek to ensure visual coherence.
-(1) A blue ogre (2) Shrek is blue 3